### v1

Abstract
Quantum many-body systems present substantial technical challenges from both analytical and numerical perspectives. Despite these difficulties, some progress has been made, including studies of interacting atomic gases and interacting quantum spins. Furthermore, the potential for criticality to enhance engine performance has been demonstrated, suggesting a promising direction for future investigation. Here, we explore the performance of a quantum Otto cycle using a long-range Ising chain as the working substance. We consider an idealized cycle consisting of two adiabatic transformations and two perfect thermalizations, eliminating dissipation. Analyzing both engine and refrigerator modes, we investigate the influence of particle number, varied from $10$ to $100$, on efficiencies and behavior near the critical point of the phase transition, which we characterize using a scaling factor. We also examine how internal factors—specifically, the power-law exponent, the number of particles, and the hot and cold reservoir temperatures—affect the system’s operation in different modes. Our results reveal that these factors have a different impact compared to their classical counterparts.
I Introduction
Thermodynamics is a remarkable theory in physics, primarily aimed at studying heat [1]. It has been instrumental in the design of internal combustion engines and even in explaining the behavior of black holes. Technological progress increasingly requires the miniaturization of components [2], pushing them to a regime where quantum effects are no longer negligible. The necessity of understanding the underlying mechanisms of converting heat into useful work has led researchers to formulate the three laws of thermodynamics. More recently, with the rise of quantum technologies and the miniaturization of devices that exchange heat and work at the nanoscale, it has become increasingly relevant to investigate these mechanisms within the framework of quantum mechanics [3]. However, while classical thermodynamics is a well-established theory, its extension to the quantum domain presents conceptual challenges and remains an active area of research. Pioneering efforts to generalize the concepts of heat and work for quantum systems date back to the 1980s [4]. Only recently has a renewed interest in quantum thermal machines spurred prolific scientific advancements, starting with the groundbreaking proposal of the maser [5] as the first example of a quantum machine.
A quantum thermal machine is broadly defined as a device made of quantum systems capable of performing work by undergoing a suitable thermodynamic cycle. Several studies have proposed potential implementations of few-body thermal machines [6] based on the Carnot [8, 7], Otto [9, 10], Stirling [11, 12], Brayton [13], and Diesel cycles [14]. Some of these realizations have been achieved using various quantum systems, including infinite potential wells [15], nuclear magnetic resonance (NMR) [16], ultra-cold atoms [17]. However, the role played by many-body interactions in the thermodynamic performance of quantum machines [18, 19, 20] is still not fully understood. Determining whether quantum machines with many-body interactions can outperform classical ones, and under what conditions this occurs, is currently at the forefront of debates in quantum thermodynamics. Long-range interacting systems [21, 22] offer promising prospects for quantum technological applications due to their resilience to external disturbances [23]. This stability enables control over the impact of dynamically generated excitation, thereby mitigating their detrimental effects. Specifically, the dynamical phase transition [24, 25] is clear when $\alpha<1$. But when $\alpha>1$, the critical point grows into a chaotic crossover region where the dynamics and asymptotic state are very dependent on the system’s parameters.
A promising direction for exploring quantum advantage lies in using many-body quantum systems as the working substance in thermodynamic engines [26, 27]. Although this approach is hindered by technical challenges, both analytically and numerically, some preliminary results have been obtained [28, 29, 30]. It has been shown that quantum criticality can play a crucial role in enhancing engine performance [31, 32, 33]. However, the exact role that many-body interactions play in the thermodynamic efficiency of quantum engines remains poorly understood. Besides, determining whether quantum engines featuring such interactions can outperform classical ones, and under what conditions, is a key question in the current quantum thermodynamics debate. One of the main obstacles lies in the difficulty of solving the dynamics of these many-body systems. In this paper, we contribute to this burgeoning field by presenting an in-depth study of the quantum Otto engine based on the Ising model (see Fig.1), focusing particularly on the potential influence of quantum criticality on the engine’s thermodynamic performance.
This article is structured as follows: First, we will present our system and the Otto cycle, describing its applications. We will explore the impact of internal parameters such as $N$, $h_{i}$, $\alpha$, $\beta_{c}$, and $\beta_{h}$ on the operating modes and highlight the quantum features absent in classical systems, as well as the transition between them. Next, we will look at the quantum refrigerator (or thermal engine) mode and see how the chain size (N) and thermal correlations change the quantum refrigerator capability $Q_{c}/N$ (or $W/N$) and the efficiency $\eta_{R}$ (or $\eta$). We will fix the time of adiabatic compression throughout the study, unlike in previous works. We will address the scenario of a perfect quantum thermodynamic machine and a refrigerator, both of which experience an infinitely slow thermodynamic phase. We will calculate the work per spin, the efficiency, and the influence of $N$ on the extraction of work per spin and its effect on efficiency. We will also perform the same analysis for the refrigerator and also describe the performance and the influence of perturbations and dissipation, using the scaling factor [34] $\Pi_{R}/N$ (or $\Pi/N$), which serves as a reliable indicator for stability. In the following section, we will examine the behavior of correlations [24, 25] for $\alpha$ ranging from 0.01 to 1.3 with a fixed $N=100$, focusing on quantum refrigerators and thermal engines. We will study how the efficiency behaves as a function of $Q_{c}/N$ and $h_{i}$ (or $\eta$ as a function of $W/N$ and $h_{i}$), and analyze their distribution across chains of varying sizes. Our final step is to calculate the proportionality between the scaling factor per spin for $N$ from $10$ to $100$, as well as the behavior of peaks before and after the phase transition and at the critical field.
II quantum otto cycle
In this section, we will describe the system and outline its phase transition. We will then examine the thermodynamic cycle and the associated physical quantities, such as the heat exchanged and the work done. Afterward, we will explore how the interaction range affects the efficiency of the cycle. The analysis will focus on how both short-range and long-range interactions influence the system’s performance, particularly near the critical point.
II.1 Many-body quantum Ising model
We study the one-dimensional transverse-field Ising model with long-range (LRIM) power-law decaying interactions. We consider an open boundary condition, which is described by the following Hamiltonian
$H=-\frac{1}{K(\alpha)}\sum_{i<j}^{N}\frac{J}{|i-j|^{\alpha}}\hat{s}_{i}^{x}\hat{s}_{j}^{x}-h\sum_{i=1}^{N}\hat{s}_{i}^{z},$ (1)
where $s_{i}^{\mu}$ are spin matrices acting on site $i$ and $\alpha$ is the power-law exponent. The Kac normalization constant, defined as
$K(\alpha)=\frac{1}{N-1}\sum_{i<j}\frac{1}{|i-j|^{\alpha}},$ (2)
ensures that the energy density remains intensive for $\alpha\leq 1$. For $\alpha=\infty$, this model reduces to the standard transverse field Ising model (TFIM), which can be exactly solved using the Jordan-Wigner transformation [36, 35] . The long range exhibits a quantum phase transition from the ferromagnetic to the paramagnetic phase at $h=2J$. This transition persists as $\alpha$ decreases [37], with the transition point shifting towards higher magnetic field values.
At the opposite extreme, $\alpha=0$ [38, 39]represents a fully connected regime, which can be analytically treated for its static and dynamic properties. In this case, the model shows long-range ferromagnetic order at low finite temperatures. This configuration is particularly interesting for $\alpha<2$, revealing exotic phenomena such as prethermalization [21, 40], dynamical phase transitions [25, 24], and dynamical confinement [41, 42].
This model is strongly influenced by the parameter $\alpha$. For instance, for $\alpha<2$, long-range ferromagnetic order is observed at finite temperatures [43], along with unique dynamic behaviors that have garnered significant experimental interest. Trapped ion systems, with their adjustable transverse fields and long-range interactions, provide an ideal framework for studying these phenomena. The critical magnetic field is estimated using the Cluster approach and Mean-Field Theory (CMFT) [44, 45, 46] to explore the phase diagram of an Ising spin chain with long-range interactions. This method captures short-range correlations while treating interactions between clusters at a mean-field level, revealing the emergence of a chaotic region for specific values of the parameter $\alpha$. In our study, we avoid mean-field or cluster decomposition approximations and spin interaction is uniform. Nevertheless, we observe significant concordance with the results obtained via CMFT, especially concerning static phase transitions. This agreement underscores the robustness of our method, providing an accurate description of phase transitions without resorting to simplifications. From this observation, it is possible to extract a qualitative behavior of the quantum critical point as a function of $\alpha$. Interpolating the data, we find
$h_{c}(\alpha)\sim\begin{cases}1&\text{if }\alpha\leq 1,\\ 0.35(3.2-\alpha)&\text{if }1<\alpha<2.\end{cases}$
Since our method is essentially a generalized mean field theory, we will use it when the latter describes the equilibrium transition ($\alpha<5/3$). Despite this, it must be stressed that CMFT is exact for large $\ell$. We will use a numerical diagonalization of the Hamiltonian in the disordered case similar to [10], where we will assume $J_{ij}=\frac{1}{K(\alpha)}\frac{J}{|i-j|}$, to compute $\omega_{k}(h)$ for more details consult Otto [35]. In conclusion, the transverse field Ising model with power-law decaying interactions offers a rich theoretical framework for exploring quantum phase transitions and complex dynamics while being directly applicable to concrete experimental setups.
II.2 Cycle description
It describes a one-dimensional long-range Ising system of $N$ quantum spins interacting with a ferromagnetic coupling strength $J>0$, in the presence of magnetic field $h$. Hereafter, we set $J=0,5$ as the energy scale and work in units where $\hbar=k_{B}=1$.
Figure 1: The schematic of the Ising Long-range quantum Otto engine consists of two adiabatic phases and two thermalization phases. From (A → B), the system undergoes an adiabatic process where the transverse field increases from $h_{i}$ to $h_{f}$. In (B → C), the system thermalizes at temperature $T_{h}$. Then, from (C → D), it is adiabatically brought back to $h_{i}$, followed by thermalization at temperature $T_{c}$ in (D → A), completing the cycle.
The long-range interaction between spins is dictated by the strength parameter $\alpha$ (we set $h_{i}<h_{f}$ and $h_{f}-h_{i}=0.5$). We will work on an ideal otto cycle illustrated in Fig.1 composed of two infinitely slow adiabatic phases and two thermalization phases without any dissipation. After that, we will examine the influence of the thermal correlations of the $N$ particles on the efficiency of the cycle. For the dynamic and ideal case, we have:
A → B: Adiabatic increase of the transverse field
The field $h(t)$ increases linearly from $h_{i}$ to $h_{f}$, while the working substance is kept isolated from the thermal baths. The expression for $h(t)$ is:
$h(t)=h_{i}+0.5\frac{t}{T}$
  • For the ideal case: $T\to\infty$ and $h=h_{i}$
B → C: Thermalization with the cold bath
Once the field reaches $h_{f}$, the magnetic field is fixed, and the working substance is brought into contact with the cold bath until it reaches thermal equilibrium at temperature $T_{c}$ (we will assume a perfect thermalization).
C → D: Adiabatic decrease of the transverse field
The field $h(t)$ is decreased linearly from $h_{f}$ back to $h_{i}$ at the same rate as in step A → B, while keeping the working substance isolated from the thermal baths. The expression for $h(t)$ is:
$h(t)=h_{f}-0.5\frac{t}{T}$ (3)
  • For the ideal case: $T\to\infty$ and $h=h_{f}$
D → A: Thermalization with the hot bath.
Once the field returns to $h_{i}$ is kepet fixed, and the working substance is brought into contact with the hot bath until it returns to its initial thermal state at temperature $T_{h}$.
When analyzing the case of an infinitely slow cycle, i.e., the limit $T\to\infty$, the regime is typically referred to as adiabatic. In this regime, the unitary evolution is slow enough to satisfy the adiabatic theorem, preventing transitions between the instantaneous eigenstates of the Hamiltonian. We will calculate the following quantities based on the Appendix A [47, 48, 49]. The efficiency of the cycle can thus be expressed as:
$\eta=\frac{W}{Q_{h}}.$ (4)
For an ideal Otto cycle, strictly speaking, perfect thermalization and perfect quantum adiabatic sweeps are ideally obtained in the limits $T\to\infty$, can be written as follows:
$Q_{h}=\sum_{k}\omega_{k}(h_{f})\Delta f_{k},$
$Q_{c}=-\sum_{k}\omega_{k}(h_{i})\Delta f_{k},$
$W=\sum_{k}\left[\omega_{k}(h_{f})-\omega_{k}(h_{i})\right]\Delta f_{k},$
$\eta=\frac{W}{Q_{h}}.$ (5)
To calculate the efficiency $\eta_{R}$ of the refrigerator, defined as
$\eta_{R}=\frac{Q_{c}}{|W|}.$ (6)
where
$\Delta f_{k}\equiv f(\beta_{h},\omega_{k}(h_{f}))-f(\beta_{c},\omega_{k}(h_{i})).$ (7)
Note that, since after a single ideal cycle the system returns to the same initial state, we have $\Delta E=0$ and thus $W=Q_{h}+Q_{c}$ follows from the first law of thermodynamics.
II.3 Operations modes
The nature of our quantum heat engine is dictated by the sign of the following quantities $Q_{c}$, $Q_{h}$, and $W$. As illustrated in Fig.(2), our engine can operate in one of the following four modes:
  • E
    (Heat engine): $Q_{c}\geq 0,Q_{h}\leq 0,W\geq 0$, the engine produces work by absorbing heat from the hot reservoir.
  • R
    ( Refrigerator): $Q_{c}\leq 0,Q_{h}\geq 0,W\leq 0$, the engine removes heat from the cold reservoir and expels it to the hot reservoir by converting the absorbed work.
  • A
    (Accelerator): $Q_{c}\geq 0,Q_{h}\leq 0,W\leq 0$, the engine absorbs energy and heats the cold reservoir from the hot reservoir.
  • H
    (Heater): $Q_{c}\leq 0,Q_{h}\leq 0,W\leq 0$, the engine absorbs heat and heats both reservoirs.
Figure 2: The color code represents the following modes: green for accelerator, purple for heater, red for heat engine, and blue for refrigerator.
Figure 3: The different operating regimes of the quantum Otto engine in the $\alpha,\beta_{h}/\beta_{c},N-h_{i}$ plane. The color code represents the following modes: accelerator (green), heater (purple), heat engine (red), and refrigerator (blue).
The purpose of this section is to show the switching between operational modes and their dependencies on the number of particles $N$, the interaction strength $\alpha$ , and the ratio $\beta_{h}/\beta_{c}$. All $\alpha$ values can be considered as long-range correlations, but for comparison purposes, we will divide them into three categories: long-range, intermediate, and short-range. This is because, as we increase $\alpha$, the effect of long-range correlations decreases and weakens. We have chosen specific values of $\alpha$, $\beta_{h}$, and $\beta_{c}$ where a diversity of modes can be observed in the ideal case. Considering the specific values imposed on our cycle, we observe the following behaviors:
For Fig.3($a$), where the system size is fixed at $N=100$, with $\beta_{h}=1.76$, $\beta_{c}=2.65$, and where $\alpha$ varies from 0 to 0.6, we observe different operational modes of the quantum engine as a function of the initial magnetic field $h_{i}$. For low values of $\alpha$ (between 0 and 0.2), long-range correlations between particles impose a refrigerator mode (blue) before the critical field $h=1$. For intermediate fields ($h_{i}=1.2$ to $1.7$), the accelerator (green) and heater (purple) modes appear; we see that they shift near the critical point to higher fields, playing the role of transient modes for the engine mode (red), which is predominant at high fields ($h_{i}=1.5$ to $2.5$). For intermediate $\alpha$ values (between $0.2$ and $0.4$), we observe a significant decrease in the refrigerator mode, indicating that medium-range correlations favor the appearance of the engine mode on both sides of the critical point ($h_{i}=0.5$ to $1.5$). The transient modes shift from $h_{i}=1.2$ at $\alpha=0.2$ to $h_{i}=0.2$ at $\alpha=0.4$. At intermediate fields, the heater and accelerator modes dominate, extending to high fields, with a significant decrease in the engine mode and the emergence of a refrigerator mode at high fields. For short-range correlations ($\alpha$ between $0.4$ and $0.6$), there is a slight decrease in the refrigerator mode and the disappearance of the transient modes A (accelerator) and H (heater), while the engine mode increases as $\alpha$ goes from $0.4$ to $0.6$. The H, A, and R (refrigerator) modes shift toward higher fields. The figure illustrates the impact of the magnetic field and correlation range on the operational modes.
In Fig.3($b$) the operational behavior is studied as a function of the ratio $\beta_{h}/\beta_{c}$, ranging from $0.45$ to $0.75$, with $\beta_{c}$ varying between 2.345 and 3.911, for a system size of $N=100$, with $\beta_{h}=1.76$ and $\alpha=0.25$. For higher values of $\beta_{h}/\beta_{c}$ (between $0.65$ and $0.75$), the refrigerator mode (blue) dominates and extends around the critical field at $h_{i}\approx 1.0$. The accelerator (A) and heater (H) modes also appear, acting as transitional modes between lower fields and the heat engine (E) in regions of more intense fields. As the $\beta_{h}/\beta_{c}$ ratio decreases (between 0.65 and 0.55), the refrigerator mode shows a significant reduction, while the A and H modes expand around the critical field, taking over from the high-field regions. Meanwhile, the zone corresponding to the engine mode (E) grows in the high-field regions and, to a lesser extent, in pre-transition phase chains. When the $\beta_{h}/\beta_{c}$ ratio is particularly low (between 0.45 and 0.55), the refrigerator mode rapidly declines for low fields, while the heater mode almost completely disappears. The accelerator mode, on the other hand, emerges after the phase transition, around $h_{i}\approx 1.5$, while the engine mode (E) strengthens, covering both low and high fields. These observations indicate that as the $\beta_{h}/\beta_{c}$ ratio approaches 1, the system tends to operate in refrigerator mode. However, as this ratio decreases, there is a gradual transition to heat engine mode, marked by the appearance and reinforcement of the accelerator and heater modes.
Fig.3($c$) examines the operational zones as a function of the system size $N$, which varies from $10$ to $100$, with $\alpha=0.25$, $\beta_{h}=1.76$, and $\beta_{c}=2.65$. For small system sizes (from $10$ to $40$), the refrigerator mode (blue) shows a dominance of the engine mode (red) for low fields before the critical point, with the appearance of the refrigerator mode at $N=20$ for low $h_{i}$. For intermediate fields, the accelerator (green) and heater (purple) modes appear (from $h_{i}=1.5$ to $2$), and the system operates exclusively in the refrigerator mode for strong fields. As we move to medium-length chains (from $40$ to $70$), there is a slight increase in the refrigerator mode for low $h_{i}$, with the appearance of the accelerator and heater modes and a slight retreat of the engine mode around the critical field. We also observe a continuous migration of the accelerator and heater modes from pre-critical fields toward the critical field, favoring the refrigerator mode, which appears in the intermediate and strong fields. For long chains (from $70$ to $100$), there is an increase in the refrigerator mode, which extends from low $h_{i}$ up to the critical field, with the disappearance of the engine mode after $N=92$ for low to intermediate fields. Additionally, the heater and accelerator modes expand from intermediate to strong fields, while the refrigerator mode barely appears and disappears in strong magnetic fields, favoring the engine, heater, and accelerator modes. We observe that fixing the range of interactions and the reservoir temperatures does not stabilize the operational modes since they also depend on the chain size.
This analysis reveals the intricate interplay between system size, interaction strength, and temperature ratio $\beta_{h}/\beta_{c}$ in determining the operational modes of a quantum engine. The results demonstrate that the operational modes are influenced not only by the chosen values for $\alpha$, $\beta_{h}$, and $\beta_{c}$, but also by the system size $N$. Through the analysis of the figures, it is evident that long-range correlations favor the refrigerator mode at low values of $\alpha$, while at intermediate values, the accelerator and heater modes become predominant. Additionally, the transient modes vary depending on the applied magnetic fields, illustrating complexdynamics when crossing the critical point. It has also been shown that varying the ratio $\beta_{h}/\beta_{c}$ significantly impacts the operational modes, favoring the refrigerator mode at higher values and leading to a transition to the heat engine mode at lower values. These results underscore the importance of a thorough understanding of interactions and correlations within the system to optimize the performance of quantum engines. This study paves the way for further explorations in optimization of the parameters to effectively exploit the different operational modes for practical applications in the field of quantum energy.
II.4 Long Range Impact
Fig.4($a$) presents an analysis of the operational modes of a quantum thermal engine and a refrigerator in the ideal case. Initially, with $N$ fixed at $100$ and $\alpha$ varied from $0.01$ to $1.3$, we investigate the influence of interaction range and correlations on the efficiency and work output per particle $W/N$ of the heat engine, and on the efficacy and extracted heat per particle $Q_{c}/N$ of the refrigerator. Subsequently, we analyze the efficiency ($\eta$, $\eta_{R}$) as a function of $W/N$ and $Q_{c}/N$, respectively, for both the heat engine and refrigerator, considering system sizes $N$ ranging from $10$ to $100$. We analyze the different operational modes of a quantum heat engine and refrigerator in the ideal case; For the refrigerator, we consider $\beta_{h}=1.4$ and $\beta_{c}=1.5$ ($\beta_{h}=0.58$ and $\beta_{c}=0.6$), while for the thermal machines, $\beta_{h}=0.2$ and $\beta_{c}=0.9$ ($\beta_{h}=0.2$ and $\beta_{c}=0.667$).
Figure 4: The influence of long-range interactions on the work and the quantum refrigerator’s performance per spin as a function of $h_{i}$. Here $N$ is fixed at 100, while $\alpha$ is varied from 00 to $1.3$.
Figure 5: In the figure, $N$ is fixed at 100, and $\alpha$ is varied from 0 to 1.3 to observe the role of long-range interactions on the efficiency $\eta$ of the quantum heat engine and the coefficient of performance $\eta_{R}$ of the refrigerator as a function of $h_{i}$.
The ideal thermal machine is presented in the two figures, Fig.4($a$-$b$) and Fig.5($a$-$b$), where $W/N$ and $W/Q_{h}$ are again plotted as a function of $h_{i}$ for different ranges of $\alpha$. In this idealized scenario, we observe that for reservoirs with a large difference in reservoir temperatures (with $\beta_{h}=0.2$ and $\beta_{c}=0.667$), higher values of $\alpha$, which favor shorter correlations, result in higher work per spin $W/N$ compared to smaller and intermediate $\alpha$ before the critical point ($0.66$ for $\alpha=1.0$ to $0.77$ for $\alpha=1.3$).
There is a significant increase in $W/N$ for larger $h_{i}$, and near the critical field, a saturation of $W/N$ values occurs for larger $\alpha$, showing similar behavior. In the post-critical field region, an inversion occurs, with lower $\alpha$ ranges showing better $W/N$ than intermediate and higher ranges. This indicates that post-critical fields tend to favor dynamics that are more advantageous for short-range correlations. When the thermal gap is reduced (with $\beta_{h}=0.2$ and $\beta_{c}=0.9$), we observe higher work per spin for long-range correlations at very low magnetic fields, with a transition occurring before the phase transition, for short-range correlations, the work per spin $W/N$ improves, with a peak observed around $h_{i}=1$ at phase transition for long range correlations. The efficiency $\eta$ exhibits similar behavior, independent of the thermal gap. It is higher for long-range correlations, with a pronounced peak at low $h_{i}$. The thermal gap slightly shifts the peak but affects all correlations in the same way (see Fig.5($a$-$b$)). Let us now move on to the same study for the refrigerator in the case where the reservoires temperature difference is shorter, with $\beta_{h}=1.4$ and $\beta_{c}=1.5$, a similar pattern is observed for $\frac{Q_{c}}{N}$ , with greater work per spin for long-range correlations. The distinction becomes more evident in the efficiency $\eta_{R}$, which shows a marked rise for all $\alpha$ values, especially the higher ones. A notable saturation occurs at the critical point, causing a disruption and a decline in short-range correlations at $h_{i}=1.5$, rendering them less effective than long-range ones. At the same time, intermediate $\alpha$ values stabilize, experiencing a slight decrease but still outperforming the higher and short $\alpha$ values (red, blue).
In the case of the ideal refrigerator, Fig.4($c$-$d$) and Fig.5($c$-$d$), the first two figures depict the behavior where the ratio $\eta_{R}=\frac{Q_{c}}{|W|}$ and the coefficient of performance $\frac{Q_{c}}{N}$ are plotted against $h_{i}$ for different values of $\alpha$, with two very close reservoirs ($\beta_{h}=0.58$ and $\beta_{c}=0.6$). We observe that for low $h_{i}$ values (before $h_{i}=0.3$), short- and intermediate-range correlations are favored, allowing more heat to be extracted from the cold reservoir per spin compared to smaller $\alpha$. For slightly higher magnetic fields (from $h_{i}=0.3$ to $h_{i}=1$), there is a shift, and long-range correlations (small $\alpha$) begin to dominate. These long-range correlations saturate particularly near the critical field. After the phase transition, a significant increase in $\eta_{R}$ and $\frac{Q_{c}}{N}$ is observed for all $\alpha$ values, without a reversal. Starting from $h_{i}=1.5$, stabilization occurs, with a slight decrease in performance for large and intermediate $\alpha$, while small $\alpha$ maintains its effectiveness, indicating that long-range correlations are more favorable in the post-critical region.
In an ideal refrigerator, we also observe that for small values of $\alpha$ (e.g., from 0.1 to 0.4), both $\frac{Q_{c}}{|W|}$ and $\frac{Q_{c}}{N}$ gradually increase with $h_{i}$, reaching a peak for higher values of $\alpha$. As $\alpha$ increases, the system is able to maintain its cooling efficiency over a wider range of $h_{i}$. This suggests that, under ideal conditions, the refrigerator can sustain a stable cooling effect even as the range of interactions extends, with improved performance observed for higher values of $\alpha$.
Figure 6: Efficient $\eta$ (COP $\eta_{R}$) as a function of $h_{i}$ and W/N ($Q_{c}$/N).
In Fig.6($c$-$d$), which illustrates the effective efficiency $\eta_{R}$ as a function of $Q_{c}/N$ and the initial transverse field $h_{i}$, small values of $N$ (around 10-30) exhibit a high sensitivity to variations in $\frac{Q_{c}}{N}$ and $h_{i}$. The effective efficiency $\eta_{R}$ reaches elevated values within a narrow range of $Q_{c}/N$, indicating an unstable behavior of the system for these smaller sizes. As $N$ increases, the curves become smoother, suggesting a stabilization of the efficiency and a reduction in sensitivity to fluctuations. For higher values of $N$, the maximum value of $\eta_{R}$ shifts towards lower $Q_{c}/N$, indicating that optimal efficiency is achieved with lower heat extraction per cycle, likely due to a more effective management of interactions in a larger system. We observe a sequential behavior: before the phase transition, we have higher efficiency $\eta_{R}$ for short chains and larger heat extraction $Q_{c}/N$ from the cold reservoir for long chains. At the critical field $h_{i}=1$, a shift occurs, and the behavior changes. Long chains become more efficient, with $\eta_{R}$ exceeding that of short chains, and $Q_{c}/N$ becomes significant for small $N$, indicating an $N$-dependence of the efficiency after the phase transition. For the second case with $\alpha=1.2$, we observe a distinct behavior and greater stability compared to the previous case. This indicates that as $\alpha$ increases, short-range correlations are favored, making them less sensitive to thermal fluctuations. Consequently, they become less dependent on the phase transition, which primarily enhances efficiency and $Q_{c}/N$, both of which are more significant for short chains.
In Fig.7($a$-$b$) In the following, we will comment on the figure, which represents the thermodynamic efficiency $\eta$ plotted as a function of $h_{i}$ and the work per spin $\frac{W}{N}$. For small values of $N$ ($10$ to $40$), at low $h_{i}$, we observe a significant efficiency, although lower than that of intermediate and longer chains, with a higher work per spin for smaller $N$. As the magnetic field increases, there is a sharp drop for short chains near the critical field $h=1$, with a slight decrease for larger $N$, which increases fluctuations for small $N$ and quickly lowers $\frac{W}{N}$ for stronger chains. Meanwhile, for longer and intermediate chains, at the critical field, there is a drop in efficiency and a significant increase in the work per spin $\frac{W}{N}$. This indicates that the efficiency of the cycle strongly depends on the initial applied field and the size of the chain.
In this study, we have explored the operational modes of a quantum heat engine and a refrigerator, emphasizing the effects of long- and short-range correlations as governed by the parameter $\alpha$. By examining various values of $\alpha$ and different reservoir temperature configurations, we have identified distinct behaviors in both the quantum thermal machine and the refrigerator.
For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the pre-critical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region. This demonstrates that critical dynamics significantly impact the work per spin $W/N$ and the efficiency. In the case of larger reservoir temperature differences, $\alpha$ plays a crucial role in determining performance, especially near the critical point where saturation is observed.
For the refrigerator, long-range correlations are similarly favored post-critical field, particularly in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$. The proximity of the reservoirs’ temperatures reveals that long-range correlations maintain stable performance even after the critical transition, while intermediate and short-range correlations exhibit a decline in efficiency.
Overall, the analysis highlights the significant influence of correlation range and magnetic field strength $h_{i}$ on the thermodynamic properties of both systems. Long-range interactions provide distinct advantages in certain regimes. Moreover, the stabilization of these effects for larger system sizes $N$ suggests that larger chains offer better optimization of quantum thermal devices, while smaller chains are more sensitive to fluctuations, particularly near the critical field. This study provides valuable insights into optimizing quantum heat engines and refrigerators through the control of system parameters and interactions.
III Quantum heat engine and Refrigerator
In this section, we will start by building a quantum refrigerator using our system and analyzing the behavior of physical quantities near the critical field. We will also investigate the effect of neighboring particles on the stability of the cycle through the scaling factor. Next, we will conduct a similar study on a quantum heat engine.
III.1 Quantum Refrigerator
the refrigeration mode (R) is predominant when the thermal gap between the two reservoirs is small Fig.3. We will examine two values of $\alpha$: the first promotes long-range correlations ($\alpha=0.25$), and the second favors short-range correlations ($\alpha=1.2$) . We will also investigate the behavior of thermodynamic quantities such as efficacy, $\frac{Q_{c}}{N}$, and the scaling factor per spin, $\frac{\Pi}{N}$
We notice in Fig.7($a$-$b$) that $Q_{c}/N$ for both values of $\alpha$ is of the same magnitude , for $\alpha=0.25$, small chains perform slightly better at low fields . We observe that it peaks just before the critical point $h_{i}=1$ and then slightly declines before and after the phase transition. It stabilizes after $h_{i}=2$ The contour lines of $\frac{Q_{c}}{N}$ for small $h_{i}$ are very tightly packed, particularly for short chains, showing greater variability and a more significant decrease compared to longer chains, which have less dense contour lines. It is observed that the contour lines indicate a slower and more stable decrease for all $N$ after $h_{i}=1.5$. In the case of $\alpha=1.2$, long chains outperform short ones before and around the critical field. In the critical field, $h_{\text{critique}}=0.7$, we observe a pronounced drop, additionally, we note that short-range correlations lead to a steeper and more abrupt decline compared to long-range correlations.We notice that for $Q_{c}/N$, the contour lines are less tight, with significant values for large $N$. Unlike the previous case, there is a more substantial decrease for long and intermediate chains, while the decrease is slower for short chains. A stabilization occurs for all chain sizes starting from $h_{i}=1.5$.
Figure 7: Contour plot of $Q_{c}$ and $\eta_{R}$ as a function of $h_{i}$ and N.
Figure 8: COP of the refrigerator versus $h_{i}$, for different system sizes $N$ across the critical point.
Figure 9: Scaling factor per spin $\Pi_{R}/N$ as a function of $h_{i}$ for different N.
The plot of COP as a function of $h_{i}$ Fig.8($a$) for different $N$ at $\alpha=0.25$ for $\frac{Q_{c}}{|Q_{c}+Q_{h}|}$ shows better efficiency for small chains (10 to 40), with a faster growth than larger chains, particularly near the critical field $h_{i}=1$. There is a distinct difference in behavior between large and small $N$, as in the post-critical region there is a sharp drop for small chains (10 to 40), with stabilization occurring from $h_{i}=1.8$. For intermediate and large chains, we observe a slight increase before the critical magnetic field, which becomes more pronounced after the phase transition, Starting from $h_{i}=1.5$, a shift in efficiency is observed: larger chains exhibit greater efficiency compared to intermediate and shorter chains. This highlights a stronger size dependence after the phase transition, particularly for higher values of $h_{i}$. Beyond the critical point, small chains struggle to recover their refrigeration efficiency, with $Q_{c}/N$ remaining low and showing no significant improvement as $h_{i}$ increases. This highlights the difficulty for smaller systems to adapt to the post-transition regime, resulting in a diminished capacity to extract heat effectively. In contrast, larger chains demonstrate a partial recovery in $Q_{c}/N$ as $h_{i}$ exceeds 1.5, suggesting their greater adaptability and resilience in the altered dynamics of the post-critical regime. While they do not reach the efficiency levels observed before the transition, large systems still maintain some level of refrigeration capability, indicating their robustness in maintaining performance despite the phase change.
The contour lines Fig.7($c$) show that they are low for small $h_{i}$, particularly for long chains. As $h_{i}$ increases (with intermediate magnetic fields $h_{i}$ from 0.5 to 1.5), the lines become concave and expand as $N$ decreases, centered around the critical field where they reach a maximum for short chains before dropping sharply after higher $h_{i}$. It is noted that for strong magnetic fields, the efficiency of long chains increases gradually.
The figure Fig.8($b$) showing the efficiency for $\alpha=1.2$ as a function of $h_{i}$ and $N$ demonstrates that short chains are more efficient than long chains, similarly to previous cases. For small $h_{i}$, there is a slight and similar increase for both long and short chains. As the critical field $h_{i}=0.7$ is approached, the slope increases, more pronounced for long chains, which rise rapidly to approach the efficiency of short chains for intermediate fields. For higher magnetic fields, there is a slower increase for long chains, with stabilization around $h_{i}=2$, and a faster increase for short chains. For large chains, saturation is observed for all values of $h_{i}$, and the increase in COP is more significant for short chains, indicating that they undergo a more gradual transition and can maintain effective refrigeration for all values of $h_{i}$.
The efficiency contour lines are less tight compared to the previous case, showing an increase in efficiency as $h_{i}$ rises for all $N$, with a leveling of the efficiency near the critical field $h_{i}=0.7$. Efficiency increases in the intermediate field range, and the contour lines become vertical. For stronger fields, starting from $h_{i}=1.7$, the lines curve and become more concave, less uniform, and more efficient for short chains than for long ones, with a peak for short chains ($N=10$ to $20$) at $h_{i}=2.5$ (see Fig.7($d$)).
The scaling factor per spin, $\Pi_{R}/N$, is influenced by the parameter $\alpha$, unlike $Q_{c}/N$, where the values are not of the same order. An increase in $\alpha$ amplifies the fluctuations in our quantum refrigerator. It is also important to note that increasing $\alpha$ brings the temperatures of the reservoirs closer together. For $\alpha=0.25$, there is a pronounced peak at low $h_{i}$ values, especially for very long chains, which diminishes as $h_{i}$ increases. This decline is more significant for shorter chains near the phase transition at $h_{i}=1$, where the slope decreases and stabilizes performance for larger $h_{i}$ values. At $h_{i}=2.25$, the performance of long and short chains converges, demonstrating a lack of size dependence for higher $h_{i}$ values.For $\alpha=1.2$, before the paramagnetic phase, performance stabilizes with a slight decline as $h_{i}$ increases across all chains. At the critical field $h_{i}=0.7$, there is a sharp drop in the scaling factor, signaling underperformance and instability, particularly for longer chains, which remain more efficient than smaller $N$ but exhibit a steeper decline. As $h_{i}$ increases, there is a modest decrease followed by stabilization (see Fig.(9)).
Overall, the system’s behavior strongly depends on the chain size and the correlation regime, with each configuration offering specific advantages depending on the magnetic field. The performance per spin is better in the ferromagnetic phase, with efficiency varying depending on the type of correlation $\alpha$.
III.2 Quantum Heat Engine
We will see in this section the evolution of the work per spin ($W/N$) and efficiency ($\eta$) as a function of the initial transverse field ($h_{i}$) for different values of $\alpha$, $\beta_{h}$, and $\beta_{c}$, analyzing the effects of the phase transition and the system’s behavior for small and large chains ($N$). Each graph in Fig.(10) highlights how interactions and thermal conditions influence the system’s performance.
Figure 10: Work per spin $W/N$ as a function of $h_{i}$ for different N.
In Fig.10($a$), with $\alpha=0.9$, $\beta_{h}=0.5$, and $\beta_{c}=1.0$, it is observed that for large values of $N$ (e.g., $N>70$), the work per spin is higher in the region where $h_{i}\leq 1.0$. Beyond this point, the slope decreases, leading to a stabilization of the curve. The maximum work per spin, $W/N$, occurs around $h_{i}=0.25$, indicating a more effective capture of short-range correlations near the critical point. It is noteworthy that, across all regions, long chains exhibit a higher $W/N$ compared to short chains. After the critical point, $W/N$ declines rapidly, especially for large systems, signaling a significant shift associated with the phase transition. In contrast, for smaller systems, the maximum is less pronounced, and the work is lower, highlighting the dominance of thermal fluctuations and local correlations. When comparing the sizes of the chains, $W/N$ remains higher for long chains throughout the $h_{i}$ range from 0 to 2.5, demonstrating greater efficiency in capturing short-range correlations for longer chains. In Fig.10($b$), corresponding to $\alpha=1.2$, $\beta_{h}=0.2$, and $\beta_{c}=0.9$, the work per spin is similarly elevated for larger values of $N$, peaking around $h_{i}=0.35$, which aligns with the estimated critical point for $\alpha>1$ at $h=0.7$. Long chains are more adept at capturing short-range correlations than their shorter counterparts prior to the transition, thereby accounting for a higher $W/N$. After $h_{i}=0.7$, $W/N$ experiences a steep decline, occurring more rapidly than in the previous case, while the work per spin of long chains remains superior to that of short chains. This observation indicates that longer chains continue to effectively capture short-range correlations even subsequent to the phase transition. In contrast, shorter chains exhibit a more stable behavior, suggesting reduced sensitivity to fluctuations in $h_{i}$. The Fig.10($c$), with $\alpha=0.55$, $\beta_{h}=0.2$, and $\beta_{c}=0.667$, shorter chains ($N<30$) exhibit a higher $W/N$ for $h_{i}$ ranging from 0 up to the critical point at $h=1.0$, where a transition occurs, allowing long chains to become more efficient following the phase transition; this alteration does not apply to short-range correlations. The peak of $W/N$ is observed around $h_{i}=0.25$, indicating that smaller chains are more effective in capturing intermediate-range correlations prior to the phase transition. However, after $h_{i}=1$, longer chains ($N>70$) surpass shorter chains in terms of $W/N$, reflecting a superior ability to capture long-range correlations. This implies that for $\alpha=0.55$, the phase transition significantly alters the behavior of long chains post $h_{i}=1$, while shorter chains maintain their efficiency before this transition. Fig.10($d$), with $\alpha=0.2$, $\beta_{h}=0.6$, and $\beta_{c}=2.0$, exhibits behavior akin to the previous case, with the performance of long chains surpassing that of shorter chains before the critical point at $h=1$, particularly near $h=0.4$. For shorter chains ($N<30$), $W/N$ is elevated for $h_{i}\leq 0.4$, peaking around $h_{i}=0.8$ for longer chains. Beyond this threshold, $W/N$ declines rapidly, indicating a loss of efficiency in the smaller chains. In contrast, after the critical field, starting from $h_{i}=1$, there is a significant drop in $W/N$. This illustrates that long-range correlations in long chains at the phase transition accelerate the decline of $W/N$.
The contour lines of the work per spin $W/N$ presented in Fig.11 show behavior that depends on $\alpha$ and the thermal gap. When $\alpha$ is small and favors long-range correlations, there are two peaks: one for short chains and the other for long chains unlike $\alpha=0.25$, we observe a peak before the transition (for small $N$) and another at the critical field (for large $N$). When $\alpha=0.55$, there is a peak before the critical field and another after the phase transition, which is farther apart with a lower $W/N$ at the critical field compared to the previous case. When short-range correlations are favored ($\alpha=0.9$ and $1.2$), we observe a single peak before the critical field for long chains. The only difference between the two cases lies in the variability and a smaller decrease in $W/N$ before the phase transition for $\alpha=1.2$, with more vertical contour lines. This detailed analysis presents the behavior of the efficiency $W/Q_{h}$ and the work per spin $W/N$ as a function of $h_{i}$ for different values of $\alpha$, $\beta_{h}$, and $\beta_{c}$.
Figure 11: Contour plot of Work per spin W/N and efficient $\eta$ as a function of $h_{i}$ and N.
With $\alpha=0.9$, $\beta_{h}=0.5$, and $\beta_{c}=1.0$, the efficiency $W/Q_{h}$ shows a gradual decline as $h_{i}$ increases. The phase transition, occurring at $h=1$, corresponds to the critical point $h_{c}$ for $\alpha\leq 1$. Prior to this transition, longer chains (with lower $N$) display greater efficiency compared to shorter chains. However, near the critical point, the efficiency of longer chains aligns closely with that of shorter ones. Beyond $h=1.25$, the disparity between short and long chains becomes more apparent, accompanied by a marked decrease in efficiency for both. In the second figure, with $\alpha=1.2$, $\beta_{h}=0.2$, and $\beta_{c}=0.9$, we observe a pattern reminiscent of the case with $\alpha=0.9$, where short-range correlations are predominant. The critical field is located at $h_{i}=0.7$. Prior to the phase transition, there is a slight reduction in efficiency for chains of all sizes, with large $N$ chains exhibiting greater efficiency. At the critical point, a distinctive behavior emerges with a sharp drop in efficiency, resulting in a significant decline across all chains, leading to an identical efficiency independent of chain length in the post-critical region. Efficiency remains relatively high for small values of $h_{i}$ (between 0 and 0.5), reaching a peak at very low fields. Large chains (with higher $N$) demonstrate superior efficiency compared to smaller chains across the entire range of $h_{i}$, both before and after the critical field. This stability in larger chains following the transition highlights a regime where short-range correlations prevail, maintaining efficiency despite variations in the transverse field $h_{i}$. In the third figure, with $\alpha=0.55$, $\beta_{h}=0.2$, and $\beta_{c}=0.667$, which favors intermediate range correlations, the efficiency $W/Q_{h}$ exhibits a gradual decline as $h_{i}$ increases, with nearly identical efficiency for large $N$ values, indicating pre-saturation for long chains in the range $h_{i}=0$ to 0.5, just before the phase transition. Around the critical field $h_{c}=1$ for $\alpha\leq 1$, a stabilization is noticeable, where efficiency remains almost unchanged. Beyond $h_{i}=1.5$, a slight reduction in efficiency is observed for all chain lengths. The curve’s behavior suggests heightened sensitivity of the system to changes in $h_{i}$ near the critical point. The Fig.12($a$), with $\alpha=0.2$, $\beta_{h}=0.6$, and $\beta_{c}=2.0$, the efficiency is initially high but steadily declines as $h_{i}$ increases, reaching the critical point at $h_{i}=1$. This behavior reflects a phase transition, where the system’s sensitivity to the transverse field $h_{i}$ is notably heightened for small values of $\alpha$. The slope increases and the decline becomes more rapid for larger $h_{i}$, emphasizing the influence of long-range correlations in stronger fields. Post-transition, shorter chains demonstrate greater stability with fewer fluctuations compared to longer chains. For large $N$, the efficiency becomes more uniform, exhibiting faster saturation, more pronounced than in the case of $\alpha=0.55$, though this saturation occurs in the post-transition region, accentuated by long-range correlation effects. In stronger fields, such as $h_{i}=2$, intermediate-length chains show superior efficiency, indicating a shift in the system’s operational regime where long-range correlations become less effective for longer chains (see Fig.12).
Figure 12: Efficiency $\eta$ as a function of $h_{i}$ for different $N$.
The contour lines of efficiency $\eta$ show a dependence on $\alpha$. Similar to $W/N$, there are two distinct behaviors. For small $\alpha$, we observe a peak for long chains at very low $h_{i}$, with the contour lines becoming more horizontal for medium and large $N$, then more vertical for long chains, which is less noticeable for $\alpha=0.55$. For larger $\alpha$, there is a clear peak for medium and large chains at low $h_{i}$, with vertical contour lines where efficiency increases horizontally. This effect is more pronounced for $\alpha=1.2$, which shows slower variability compared to $\alpha=0.9$ (Fig.11(e)-(h)).
Unlike work, the efficiency for all the parameters previously mentioned is dependent on $N$. To increase efficiency for fixed parameters, it is sufficient to increase the number of particles, which is not always the case for $W/N$. We observe that the efficiency ($\eta$) and the work per spin ($W/N$) exhibit different behaviors, as the work per spin is not proportional to the chain size for small and intermediate values of $\alpha$. In contrast, the efficiency depends directly on the size of the chain: the larger $N$ becomes, the greater the efficiency ($\eta$). This discrepancy can be explained by the nature of the correlations and fluctuations induced by the temperature difference between the reservoirs. For larger values of $\alpha$ (e.g., 0.9 and 1.2), we observe that when the temperature difference ($\Delta T$) is significant (as with $\alpha=1.2$), the system achieves a higher work per spin ($W/N$) compared to $\alpha=0.9$, but with a lower efficiency. This can be attributed to thermal fluctuations and agitation, which are reduced when $\Delta T$ is smaller. The same interpretation applies to long-range and intermediate-range correlations, where the effect of temperature fluctuations diminishes with a smaller temperature gradient, leading to different performances in terms of both work and efficiency. Thus, the relationship between efficiency and work per spin reflects the complex interplay of thermal fluctuations and correlations in the system, especially across different interaction ranges.
III.3 Scaling factor of the quatum heat engine
The scaling factor $\Pi=\frac{W}{\eta_{\text{Carnot}}-\eta}$ measures how effectively an engine produces work despite having an efficiency lower than the Carnot limit. A high scaling factor means that the system is optimized to extract significant work, even in non-ideal conditions. In quantum engines, a higher $\Pi$ reflects more efficient energy conversion through quantum effects like correlations and superposition. It also suggests lower energy losses, allowing the system to approach Carnot efficiency. Overall, a high scaling factor indicates that the system is productive, efficient, and resilient to fluctuations.
To observe the impact and influence of correlations without considering the size of the chain, we will normalize by dividing by $N$.We will discussed this in more detail in the appendix Appendix D. The corrected interpretation of the provided figures focuses on the scaling factor $\Pi$, defined as
$\Pi/N=\frac{W}{(\eta_{\text{Carnot}}-\frac{W}{Q_{h}})}/N,$ (8)
where $W$ is the work done by the system, $\eta_{\text{Carnot}}=1-\frac{\beta_{h}}{\beta_{c}}$ is the Carnot efficiency, and $N$ is the number of particles in the system . The figures 13 illustrate $\Pi/N$ as a function of the transverse magnetic field $h_{i}$, highlighting the behavior of the system for varying values of $N$, ranging from 10 to 100, for different configurations of $\alpha$, $\beta_{h}$, and $\beta_{c}$. Fig.13($a$) $\alpha=0.9$, $\beta_{h}=0.5$ and $\beta_{c}=1.0$, $\Pi/N$ starts at high values for small $h_{i}$ and decreases significantly near the critical magnetic field $h_{i}=1$, indicating a strong phase transition. Large chains ($N>70$) show a sharp decrease in $\Pi/N$, while small chains show a more gradual decline. After $h_{i}=1$, $\Pi/N$ for large chains remains stable but lower compared to small $h_{i}$ values, indicating a more consistent capture of long-range effects. $\Pi/N$ drops significantly after the transition, reflecting their inability to maintain high performance in the long-range regime. In , corresponding to $\alpha=1.2$, Fig.13($b$)$\beta_{h}=0.2$ and $\beta_{c}=0.9$, $\Pi/N$ is high for large values of $N$ (for example, $N>70$) in the region where $h_{i}\leq 0.7$. The maximum occurs around $h_{i}=0.3$, reflecting better adaptation to critical fluctuations near the estimated transition point for $\alpha>1$. Beyond the critical point $h_{i}=0.7$, there is a sharp drop in the system’s performance, with a less steep slope for small chains. After $h_{i}=1$, $\Pi/N$ decreases for large chains but remains higher compared to small chains ($N<30$), indicating that large chains are more effective in maintaining high performance before the transition but experience a more intense drop in performance than small chains, where long-range correlations dominate. Large chains show a rapid decrease in $\Pi/N$, reflecting their limited capacity to harness critical effects. At $\alpha=0.55$ Fig.13($c$), $\beta_{h}=0.2$ and $\beta_{c}=0.667$, shows that $\Pi/N$ is initially high, especially for small chains. Where long-range correlations are moderate before the critical field $h=1$, in the ferromagnetic phase, they force the particles to align better than in long chains, thus ensuring better $\Pi/N$ performance. After the transition phase $h_{i}=1$, large chains dominate in terms of $\Pi/N$, reflecting their superior performance in exploiting intermediate correlations in the ferromagnetic phase at higher field strength. This behavior indicates that large chains are more adaptable and efficient after the transition, whereas small chains are more efficient in the ferromagnetic phase, before the phase transition. In Fig.13($d$), with $\alpha=0.2$, $\beta_{h}=0.6$ and $\beta_{c}=2.0$, where the interactions are the longest, we observe that before $h_{i}=0.25$, the short chains perform better, and there is a shift at this point towards the dominance of long chains. $\Pi/N$ peaks at $h_{i}=0.6$ before the critical point but decreases rapidly beyond this point, showing poor performance in the post-transition regime. We also observe that $\Pi/N$ decreases slowly after $h_{i}=1$, indicating an effective capture of long-range correlations. This highlights the dominance of long-range effects for large chains before the phase transition. It is also important to note that, as in the previous section, there is a slight shift between long and short chains post-phase transition, as the short chains stabilize more quickly because the fluctuations and thermal perturbations, amplified by long-range correlations in the paramagnetic phase.
Figure 13: Contour plot of the scaling factor per spin $\Pi_{R}/N$ as a function of $h_{i}$ and N.
The impact of $\alpha$, $\beta_{h}$, and $\beta_{c}$ on the system is significant. For high $\alpha$ values (e.g., $\alpha=1.2$), short-range correlations dominate, leading to a smoother transition and more stable $\Pi/N$. Large chains benefit significantly from short-range interactions before $h_{c}$. For low $\alpha$ before critical field long-range effects are pronounced, leading to sharper transitions in $\Pi/N$ and small chains are more effective in these regimes, showing larger values of $\Pi/N$. A higher thermal gradient enhances the sensitivity of $\Pi/N$ to variations in $h_{i}$. For example, in the third figure, the large difference between $\beta_{c}=2.0$ and $\beta_{h}=0.6$ causes a significant shift in $\Pi/N$ post-transition.faor a small thermal gap (e.g., $\beta_{c}=1.0,\beta_{h}=0.5$ in the second figure) result in a smoother decline of $\Pi/N$, with less abrupt changes across the critical point but with a lower performance than the others.
Before the critical point ($h_{i}<1$), small chains exhibit higher $\Pi/N$, reflecting their ability to efficiently capture long-range correlation in weak field $h_{i}$ , while large chains have a lower $\Pi/N$ but show a better performance than the shorter chain for higher $h_{i}$ special in the vicinity of critical point and after phase transition, indicating the growing influence of long-range correlations after the critical point ($h_{i}>1$), large chains outperform small chains, with a higher $\Pi/N$ due to the dominance of long-range correlations. Small chains show a significant decline in $\Pi/N$, reflecting reduced efficiency in capturing long-range effects. At the critical point ($h_{i}=1$), $\Pi$ peaks for large chains as the system approaches Carnot efficiency, reflecting enhanced work output and system efficiency. For small chains, the peak is less pronounced, indicating their limited capacity to utilize critical phenomena effectively.
Refer to captionRefer to caption
; Refer to captionRefer to caption
Figure 14: The evolution of the maxima as a function of $N$ (the number of particles) for the diamagnetic peak (in red) and the paramagnetic peak (in blue) of $\Pi/N$ and $\Pi_{R}/N$, with a comparison to the critical field (in red).
For the case of short-range correlations, we observe an underperformance after the phase transition, but with a stronger dependence on system size, as longer chains prove more efficient both before and after the critical point. It is also noted that for large $\alpha$ values, short-range correlations dominate, and the system’s efficiency is enhanced by the thermal gap, which promotes short thermal correlations and better protects them from thermal fluctuations.
The behavior of $\Pi/N$ as a function of $h_{i}$ demonstrates the intricate relationship between work, system efficiency, and particle number $N$. Large chains tend to dominate in regimes where long-range correlations are significant, especially beyond the critical point, whereas small chains are more effective in the short-range, pre-transition regime. The parameters $\alpha$, $\beta_{h}$, and $\beta_{c}$ play crucial roles in shaping this behavior, with their values determining the system’s sensitivity to field variations and the overall efficiency relative to Carnot. Understanding these dynamics is essential for optimizing quantum heat engines across different operating regimes.
IV Investigating the role of thermal correlations and criticality
In this section, we will examine the scaling factor $\Pi/N$, studying its impact on the system’s behavior for varying values of $N$ (ranging from 10 to 100) and different configurations of $\alpha$, $\beta_{h}$, and $\beta_{c}$. focusing on the behavior near the critical field, including the peaks observed in the paramagnetic and ferromagnetic phases. Additionally, we will discuss how these features depend on $N$ in both the quantum heat engine and refrigerator modes.
IV.1 Critical reaction
We will examine the role of the scaling factor and the behavior at the critical field for $N$ ranging from 10 to 100. We will also analyze the peaks in the paramagnetic phase before the phase transition and the ferromagnetic phase after the transition, as well as their proportionality and dependence on N for the quantum heat engine and refrigerator mode. We intend to use a regression to calculate the exponent $a$ responsible for the growth rate of the scaling factor per spin, proportional to $N^{a}$. We will then examine the behavior of the performance and its trend for small, intermediate, and long chains and the role of thermal gaps and correlations $\alpha$ , to observe the impact of the thermal reservoirs on our system. In the case where $\alpha=0.9$, with $\beta_{h}=0.5$ and $\beta_{c}=1.0$ Fig.14($a$), the obtained exponents are $a=0.4578$ for the critical field, and 0.5154 and 0.46 respectively for the ferromagnetic and paramagnetic phases , with a poor temperature difference between the reservoirs. The scaling factor per particle describes the performance per particle while accounting for the influence of others. We observe that the post-critical and critical exponents are almost identical, with a significant slope for small $N$ (10 to 40) and medium $N$ (40 to 70). This suggests that critical effects slow down the growth and stabilize it in the paramagnetic phase, especially for large $N$ (70 to 100). In the ferromagnetic phase, we see greater growth, with a slight decrease in slope for larger $N$, and the curve is higher than in the critical and paramagnetic cases. For $\alpha=1.2$, with $\beta_{h}=0.2$ and $\beta_{c}=0.9$ Fig.14($b$), the obtained exponents are $a=0.2355$ for the critical field, 0.2308 and 0.2355 for the ferromagnetic and paramagnetic phases , respectively. In this case, $\alpha>1$, which favors more short-range correlations, and the thermal gap is higher compared to the previous case. The exponents are smaller than for $\alpha=0.9$ and are very close to each other, with one of the curves (critical and post-critical) perfectly overlapping The paramagnetic phase exhibits slightly lower growth than the ferromagnetic phase, but both show faster growth with $N$. A similar behavior is observed for small to intermediate $N$, but a notable difference appears for large $N$, where the critical and paramagnetic curves show the onset of pre-saturation the slope increases only marginally between $N=90$ and $N=100$. This indicates that for larger $\alpha$, the system’s performance depends on its size, which significantly amplifies thermal fluctuations and correlations, potentially leading to a more optimal yield. For Fig.14($c$) $\alpha=0.55$, which favors intermediate correlations, with a high thermal gap $\beta_{h}=0.2$ and $\beta_{c}=0.667$ Before the critical field, we observe a curve that is higher than the critical field and the post-phase transition, but it shows a downward behavior indicating underperformance of intermediate chains, which slows down around $N=90$. Near $N=55$ In the ferromagnetic phase, the spins are better aligned, but increasing the number of particles makes the system more sensitive to fluctuations caused by the different types of correlations, which increase the noise and cause the decrease of the scaling factor per spin ($\alpha=-0.1035$). There is an overestimation of the scaling factor per spin of the post-critical field peaks ($\alpha=0.2432$), with an increase of the performance of long chains. At the critical point, there is a slight and slow increase for long chains ($\alpha=0.0504$), indicating a greater difficulty aligning the particles for intermediate correlations before and during the phase transition. In the case of $\alpha=0.2$, with $\beta_{h}=0.6$ and $\beta_{c}=2.0$, Fig.14($d$) the thermal gap is short, low values of $\alpha$ favor long-range correlations, indicating that the fields before the critical point are more efficient. For all chains, performance in the phase transition and after the critical field shows a slight decrease in intermediate N, while the curves for the critical and post-critical peaks demonstrate an increase and outperform of the high chains, which is proportional to $N$.A slight under performance is observed in long chains (from $70$ to $100$) before the phase transition. The scaling factor per spin in the critical fields and after the transition the curves are superimposed for intermediate and shorter chains. This indicates that in long correlation before the critical field there is a superior performance and better spin alignment, resulting in less dissipation and a higher efficiency, there is nonetheless slower growth ($\alpha=0.32826$) compared to postcritical fields ($\alpha=0.9434$) and at the phase transition ($\alpha=0.9413$). We will then move on to the refrigeration mode, which is more prevalent in regions with a small thermal gap. We will examine the influence of the parameter $\alpha$, considering both small and large values, which, respectively, promote long-range and short-range thermal correlations. In the case of $\alpha=0.25$, with $\beta_{h}=1.4$ and $\beta_{c}=1.5$, Fig.14($e$)there is a perfect superposition for intermediate and long chains, with a slight difference for small $N$ for the critical ($\alpha=1.6150$) and postcritical fields ($\alpha=1.2702$). We observe that the performance per particle of the post phase transition for all N is lower than the pre phase transition, which exhibits a significant increase for small chains and saturates at intermediate $N$ ($\alpha=0.5338$), then slightly decreases. In contrast, the critical field reduces the gap for larger chains. It is also noted that the performance before the phase transition is less dependent on the size, which is an effect of long-range correlations. For of $\alpha=1.2$, with $\beta_{h}=0.58$ and $\beta_{c}=0.6$, Fig.14($f$) there is a favoring of short-range correlations. Similarly, there is a perfect superposition between the critical fields and the post-critical peaks ($\alpha=0.4717$) and the the pics before the phase transition increase more slowly ($\alpha=0.4071$), however unlike the previous case they exhibit similar behavior with an increase as the chain size increases, but with a slightly smaller slope for the pre-critical case, which remains higher than the others regardless of $N$. This demonstrates that short-range correlations impose a chain-length dependency.
The scaling factor $N^{\alpha}$ is crucial for understanding how the yield of a quantum thermal engine behaves as the system size increases. The goal is to identify conditions under which increasing $N$ (number of spins or system size) leads to a significant improvement in yield compared to the energy cost. When $\alpha>0$, it means that increasing $N$ results in a superlinear growth of cycle work and engine efficiency, which is a sought after goal to achieve better performance on a large scale. Near a critical point, energy fluctuations and heat capacity can diverge, leading to faster yield growth (higher value of $\alpha$). This is where critical engines can potentially achieve efficiency close to the Carnot point without sacrificing power. For lower values of $\alpha$ (typically far from the critical point), the yield per spin reaches saturation, as spin interactions are not coherent enough to benefit from increasing system size. In conclusion, the results show that the dynamics of thermal efficiency and work per spin depend strongly on the value of $\alpha$, which defines the range of interactions between spins. Temperatures ($\beta_{h}$ and $\beta_{c}$) also influence the nature of transitions, with marked differences in spin thermal sensitivity at different values of $h_{i}$. The observed critical points are generally in agreement with theoretical predictions, although some deviations suggest significant correlation effects in these regimes.
We generally observe better performance when the thermal gaps are very large for $\alpha=1.2$ and $\alpha=0.55$. When short-range correlations are favored, we see a more linear critical behavior that depends more on $N$, particularly with the phase transition. For lower $\alpha$, there is less dependence on the chain length, which can experience a drop of $\pi/N$ for long-range drops ($\alpha=0.55$), or similar behavior before and after the phase transition for large $N$ ($\alpha=0.2$).
In conclusion, the performance of quantum thermal machines, whether as engines or refrigerators, is intricately linked to the scaling behavior governed by $\alpha$. High values of $\alpha$ in engines suggest promising avenues for enhancing efficiency through system size, particularly near critical points where spin correlations are maximized. However, refrigerators, even with intermediate values $\alpha$, show limited improvement, pointing to the need for optimizing interaction dynamics and minimizing dissipation to achieve better performance. Understanding these scaling laws and their implications can guide the design and optimization of next-generation quantum thermal devices, tailored for specific operational needs.
V Conclusion
In this study, we investigated the performance of a long-range Ising chain quantum Otto engine and refrigerator under ideal conditions, focusing on the effects of system size $N$, the interaction range parameter $\alpha$, and the temperature ratios $\beta_{h}/\beta_{c}$. Our analysis demonstrates that thermal correlations significantly influence the operational efficiency of these quantum machines, particularly near the critical points of phase transitions. We examined the behavior of all types of correlations for a long chain (N=100). For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the precritical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region. Similarly, for the refrigerator, long-range correlations are also favored in the post-critical field, especially in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$.
We then fixed $\alpha$ and varied $N$ to investigate the influence of chain size on performance. For the quantum refrigerator, our results indicate that the cooling efficiency is maximized for smaller system sizes before the phase transition. At the critical point, the behavior of the efficiency $\eta_{R}$ strongly depends on the value of $\alpha$. For long-range correlations, efficiency decreases for short chains but increases for long chains. In contrast, for short-range correlations, $\eta_{R}$ increases, showing better performance for short chains, while long chains exhibit greater stability and reduced dissipation. For the quantum heat engine, we observed relatively stable performance with minimal efficiency loss when the thermal gap is close and transitions occur before the phase transition, where long-range and intermediate correlations show better performance, even as the system size increases. This stability is attributed to the absence of non-adiabatic effects and the slow, quasi-static nature of the transformation. Additionally, the yield $\eta$ differs from that of the refrigerator, as it depends on the chain size for all values of $\alpha$.
Overall, this study highlights the importance of distinguishing between the critical phase and the diamagnetic phase in understanding the performance of quantum heat engines and refrigerators. The ferromagnetic phase, characterized by significant fluctuations and superior performance, contrasts with the diamagnetic phase, which occurs beyond the critical point and exhibits greater stability but inferior performance. These findings underscore the need for a deep understanding of both static and dynamic conditions to effectively optimize the operation of quantum thermal machines.
Appendix A Dynamic of the system
A stationary state is reached in the long-time limit $t\to\infty$. In the following, we will characterize this stationary state through observables. We start with the adjoint version of our master equation, which can be written as:
$\frac{dO_{H}(t)}{dt}=i\{H,O_{H}(t)\}_{-}+\sum_{n,k}\gamma_{n,k}\Big{(}[1-\zeta f_{n}(\omega_{k})]\\ \times[2b_{k}^{\dagger}O_{H}(t)b_{k}-\{b_{k}^{\dagger}b_{k},O_{H}(t)\}_{+}]\\ +f_{n}(\omega_{k})[2b_{k}O_{H}(t)b_{k}^{\dagger}-\{b_{k}b_{k}^{\dagger},O_{H}(t)\}_{+}]\Big{)},$ (9)
where $O_{H}(t)$ denotes the Heisenberg form of a Schrödinger observable $O$.
$\frac{d}{dt}\langle b_{k}^{\dagger}b_{k}\rangle$ $=-2\sum_{n}\gamma_{n,k}\langle b_{k}^{\dagger}b_{k}\rangle+2\sum_{n}\gamma_{n,k}f_{n}(\omega_{k}),$ (64a)
$\frac{d}{dt}\langle b_{k}^{\dagger}b_{q}\rangle$ $=\Big{(}i(\tilde{\omega}_{k}-\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{k}^{\dagger}b_{q}\rangle,$ (64b)
$\frac{d}{dt}\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle$ $=\Big{(}i(\tilde{\omega}_{k}+\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle,$ (64c)
$\frac{d}{dt}\langle b_{q}b_{k}\rangle$ $=\Big{(}-i(\tilde{\omega}_{k}+\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{q}b_{k}\rangle.$ (64d)
Developing the expressions, we find:
$\langle b_{k}^{\dagger}b_{k}\rangle(t)$ $=\frac{\sum_{n}\gamma_{n,k}f_{n}(\omega_{k})}{\sum_{n}\gamma_{n,k}}\Big{(}1-\exp\big{(}-2\sum_{n}\gamma_{n,k}t\big{)}\Big{)}$
$\quad+\langle b_{k}^{\dagger}b_{k}\rangle_{0}\exp\big{(}-2\sum_{n}\gamma_{n,k}t\big{)},$ (65a)
$\langle b_{k}^{\dagger}b_{q}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}\rangle_{0}\exp\Big{(}i(\tilde{\omega}_{k}-\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65b)
$\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{0}\exp\Big{(}i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65c)
$\langle b_{q}b_{k}\rangle(t)$ $=\langle b_{q}b_{k}\rangle_{0}\exp\Big{(}-i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65d)
where the subscript $\langle\cdot\rangle_{0}$ indicates the expectation value at time $t=0$. If $\sum_{n}\gamma_{n,k}\neq 0$, the limit $t\to\infty$ gives:
$\langle b_{k}^{\dagger}b_{q}\rangle_{s}$ $=\delta_{kq}\frac{\sum_{n}\gamma_{n,k}f_{n}(\omega_{k})}{\sum_{n}\gamma_{n,k}},\quad\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{s}=\langle b_{q}b_{k}\rangle_{s}=0.$ (66)
The equations (65) cease to be valid if, for some pair $(k,q)$, $\gamma_{n,k}=\gamma_{n,q}=0$ for all $n$. In that case, we find:
$\langle b_{k}^{\dagger}b_{q}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}\rangle_{0}e^{i(\tilde{\omega}_{k}-\tilde{\omega}_{q})t},$ (67a)
$\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{0}e^{i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t}.$ (67b)
Appendix B Adiabatic Transformation
The time integral can be evaluated (assuming $t$ is large) as:
[…]
The transition amplitude becomes:
$a_{n\to m}(t)\approx-\frac{2i}{\Delta_{mn}}\langle m(\lambda)|\frac{dH(\lambda)}{d\lambda}|n(\lambda)\rangle\frac{d\lambda}{dt}\sin\left(\frac{\Delta_{mn}t}{2\hbar}\right)$ (11)
The transition probability $P_{n\to m}$ is the square of the transition amplitude:
$P_{n\to m}(t)$ $=|a_{n\to m}(t)|^{2}\approx\left(\frac{2}{\Delta_{mn}}\right)^{2}\left|\langle m(\lambda)\middle|\frac{dH(\lambda)}{d\lambda}\middle|n(\lambda)\rangle\right|^{2}$
×$\quad\times\left(\frac{d\lambda}{dt}\right)^{2}\sin^{2}\left(\frac{\Delta_{mn}t}{2\hbar}\right)$ (12)
For small transitions, the oscillating term $\sin^{2}\left(\frac{\Delta_{mn}t}{2\hbar}\right)$, which oscillates between 0 and 1, can be ignored. Thus, the transition probability simplifies to:
$P_{n\to m}\propto\left(\frac{1}{\Delta_{mn}^{2}}\right)\left|\langle m(\lambda)|\frac{dH(\lambda)}{d\lambda}|n(\lambda)\rangle\right|^{2}\left(\frac{d\lambda}{dt}\right)^{2}$ (13)
We have demonstrated that the transition probability $P_{n\to m}$ is proportional to:
$P_{n\to m}\propto\left(\frac{1}{\Delta_{mn}^{2}}\right)\left|\langle m(\lambda)|\frac{dH(\lambda)}{d\lambda}|n(\lambda)\rangle\right|^{2}\left(\frac{d\lambda}{dt}\right)^{2}$ (14)
This result forms the basis of the adiabatic condition, which states that for a process to be adiabatic, the rate of change $\frac{d\lambda}{dt}$ must be much smaller than the square of the energy difference $\Delta_{mn}^{2}$, thereby ensuring that the transition probability remains low.
Appendix C Non perfect Thermalization
In our paper, we worked on an ideal cycle with perfect ideal and adiabatic transformations, assuming the system is in contact with the two reservoirs for a time $t<a$, where $a$ is the thermalization time. According [18], the distribution becomes:
$\langle b_{k}^{\dagger}b_{k}\rangle(\gamma)=f(\beta,\omega_{k})(1-e^{-J\gamma})+\langle b_{k}^{\dagger}b_{k}\rangle_{\rho_{0}}e^{-J\gamma},$ (15)
We have two thermalization phases, from B to C and from D to A, each lasting $a/2$, and two adiabatic phases, from A to B and from C to D, each lasting $\gamma/2$. We assume $J=1$ here. Note that $J$ is a parameter that depends on the properties of the bath and the system-bath coupling, and obviously affects the thermalization time scale. The off-diagonal correlators and anomalous terms $\langle b_{k}b_{q}\rangle$ remain zero at any time, provided the initial condition for the Otto cycle is a thermal state (say, $\rho_{c}$).
Let’s find an expression for the distribution after $n$ repetitions of the non-perfectly thermalized cycle. We introduce the diagonal $N\times N$ matrices :
$\Theta_{c}(h)=\text{diag}\left(f\left[\beta_{c}(h),\omega_{k}\left(\lambda_{i}(f)\right)\right]\right)_{k=1,\ldots,N},$ (16)
$\Gamma_{c}^{[n]}(h)=\text{diag}\left(\text{Tr}\left[b_{k}^{\dagger}b_{k}\rho_{c}^{[n]}(h)\right]\right)_{k=1,\ldots,N},$ (17)
where $\rho_{c}^{[n]}(h)$ denotes the state at the end of the non-perfectly thermalized cycle with the bath at temperature $\beta_{c}^{-1}(h)$ after the $n$-th cycle repetition.
At the end of the ideal adiabatic transformation, the system is put in contact with the reservoir at temperature $\beta^{-1}_{h}$ and reaches the state $\rho^{[1]}_{h}$, characterized by the distribution:
$\Gamma^{[1]}_{h}=\Theta_{h}(1-e^{-\gamma})+\Gamma^{[0]}_{c}e^{-\gamma}.$ (18)
After the ideal adiabatic transformation back, the system is put in contact with the reservoir at temperature $\beta^{-1}_{c}$ reaching the state $\rho^{[1]}_{c}$, characterized by the distribution:
$\Gamma^{[1]}_{c}=\Theta_{c}(1-e^{-\gamma})+\Gamma^{[1]}_{h}e^{-\gamma}.$ (19)
By imposing the stationary condition that the distributions $\Gamma$ are no longer dependent on the iteration, we obtain the following stationary solution:
$\Gamma^{\infty}_{c}=h(\gamma)\left(\Theta_{c}+e^{-\gamma}\Theta_{h}\right),$ (20)
$\Gamma^{\infty}_{h}=h(\gamma)\left(\Theta_{h}+e^{-\gamma}\Theta_{c}\right),$ (21)
where $h(\gamma)=(1+e^{-\gamma})^{-1}$.
The equations above suggest that, for the non-perfectly thermalized cycle ($n$-th), the heat absorbed and the work done per cycle are given by:
$Q^{n\text{-th}}_{h}=g(\gamma)Q_{h},\quad W^{n\text{-th}}=g(\gamma)W,$ (22)
where $g(\gamma)=(1-e^{-\gamma})h(\gamma)=\tanh(\gamma/2)$.
The efficiency of the thermal engine in the case of the non-perfectly thermalized cycle remains the same as in the ideal case.
For the case where we have a cycle with thermalization time and adiabatic time ($\tau=a/2$) and quench time ($T=b/2$):
$P(a,b)=\frac{W^{a+b}}{a+b}=p(a,b)W,$ (23)
where the function $p(a,b)$ is defined by:
$p(a,b)\equiv\frac{\tanh(b/2)}{a+b}.$ (24)
This relation shows that the power of the engine depends on the two parameters $a$ and $b$, which respectively characterize the quench time and the thermalization time.
Appendix D Scaling Factor
A high scaling factor ($\Pi$) indicates that, despite having an efficiency ($\eta$) lower than the Carnot efficiency, the engine is able to generate a significant amount of work. This suggests that the system is effectively optimized to extract as much work as possible, even under conditions far from the ideal efficiency. In quantum engines, a higher scaling factor reflects a more efficient conversion of thermal energy into work by leveraging quantum effects such as correlations and superposition. This implies that the system is able to utilize its quantum properties more effectively to enhance its thermodynamic performance. Furthermore, a high scaling factor can also indicate reduced energy losses, such as thermal dissipation or decoherence, allowing the engine to better harness the absorbed energy and maintain an efficiency closer to the Carnot limit. Thus, a higher scaling factor signifies a system that is not only more efficient but also more resilient to fluctuations and phase transitions, making better use of available energy resources to produce work, even if its absolute efficiency remains below that of Carnot. In summary, it reflects the system’s ability to maintain a favorable balance between the generated work and its deviation from the maximum theoretical efficiency.
In conclusion, a high scaling factor is a global indicator of performance within a system, suggesting not only better energy conversion into work but also increased resilience to perturbations and more efficient use of available energy. This applies to both classical and quantum engines, where the effects of fluctuations and phase transitions become significant.
ACKNOWLEDGMENTS
A.H. acknowledges the financial support of the National Center for Scientific and Technical Research (CNRST) through the “PhD-Associate Scholarship-PASS” program.

### v2

@@
-
+Multiparticle Quantum Heat Engine: Exploring the Impact of Criticality on Efficiency
 Abstract
@@
 Technological progress increasingly requires the miniaturization of components [2], pushing them to a regime where quantum effects are no longer negligible.
-The necessity of understanding the underlying mechanisms of converting heat into useful work has led researchers to formulate the three laws of thermodynamics.
-More recently, with the rise of quantum technologies and the miniaturization of devices that exchange heat and work at the nanoscale, it has become increasingly relevant to investigate these mechanisms within the framework of quantum mechanics [3].
+The necessity of understanding the underlying mechanisms of heat conversion into useful work has led researchers to formulate the three laws of thermodynamics.
+More recently, with the rise of quantum technologies and the miniaturization of devices that exchange heat and work on the nanoscale, it has become increasingly relevant to investigate these mechanisms within the framework of quantum mechanics [3].
 However, while classical thermodynamics is a well-established theory, its extension to the quantum domain presents conceptual challenges and remains an active area of research.
@@
 Some of these realizations have been achieved using various quantum systems, including infinite potential wells [20], nuclear magnetic resonance (NMR) [21], ultra-cold atoms [22].
-However, the role played by many-body interactions in the thermodynamic performance of quantum machines [18, 19, 20] is still not fully understood.
+However, the role played by many-body interactions in the thermodynamic performance of quantum machines [23, 24, 25] is not yet fully understood.
 Determining whether quantum machines with many-body interactions can outperform classical ones, and under what conditions this occurs, is currently at the forefront of debates in quantum thermodynamics.
@@
 Specifically, the dynamical phase transition [30, 31] is clear when $\alpha<1$.
-But when $\alpha>1$, the critical point grows into a chaotic crossover region where the dynamics and asymptotic state are very dependent on the system’s parameters.
-A promising direction for exploring quantum advantage lies in using many-body quantum systems as the working substance in thermodynamic engines [26, 27].
+But when $\alpha>1$, the critical point grows into a chaotic crossover region where the dynamics and asymptotic state are very dependent on the system’s parameters.A promising direction for exploring quantum advantage lies in using many-body quantum systems as the working substance in thermodynamic engines [34, 35].
 Although this approach is hindered by technical challenges, both analytically and numerically, some preliminary results have been obtained [36, 37, 38].
@@
 However, the exact role that many-body interactions play in the thermodynamic efficiency of quantum engines remains poorly understood.
-Besides, determining whether quantum engines featuring such interactions can outperform classical ones, and under what conditions, is a key question in the current quantum thermodynamics debate.
-One of the main obstacles lies in the difficulty of solving the dynamics of these many-body systems.
-In this paper, we contribute to this burgeoning field by presenting an in-depth study of the quantum Otto engine based on the Ising model (see Fig.1), focusing particularly on the potential influence of quantum criticality on the engine’s thermodynamic performance.
-This article is structured as follows: First, we will present our system and the Otto cycle, describing its applications.
-We will explore the impact of internal parameters such as $N$, $h_{i}$, $\alpha$, $\beta_{c}$, and $\beta_{h}$ on the operating modes and highlight the quantum features absent in classical systems, as well as the transition between them.
-Next, we will look at the quantum refrigerator (or thermal engine) mode and see how the chain size (N) and thermal correlations change the quantum refrigerator capability $Q_{c}/N$ (or $W/N$) and the efficiency $\eta_{R}$ (or $\eta$).
-We will fix the time of adiabatic compression throughout the study, unlike in previous works.
-We will address the scenario of a perfect quantum thermodynamic machine and a refrigerator, both of which experience an infinitely slow thermodynamic phase.
-We will calculate the work per spin, the efficiency, and the influence of $N$ on the extraction of work per spin and its effect on efficiency.
-We will also perform the same analysis for the refrigerator and also describe the performance and the influence of perturbations and dissipation, using the scaling factor [34] $\Pi_{R}/N$ (or $\Pi/N$), which serves as a reliable indicator for stability.
-In the following section, we will examine the behavior of correlations [24, 25] for $\alpha$ ranging from 0.01 to 1.3 with a fixed $N=100$, focusing on quantum refrigerators and thermal engines.
-We will study how the efficiency behaves as a function of $Q_{c}/N$ and $h_{i}$ (or $\eta$ as a function of $W/N$ and $h_{i}$), and analyze their distribution across chains of varying sizes.
-Our final step is to calculate the proportionality between the scaling factor per spin for $N$ from $10$ to $100$, as well as the behavior of peaks before and after the phase transition and at the critical field.
+The study of topological phases in systems with long-range (LR) interactions represents a dynamic research frontier, bolstered by recent experimental proposals.
+Notably, chains with non-local hopping and pairing are promising candidates for realization [42] in solid-state systems,an example of the rigidity of long-range interacting platforms against external drivings, is the possibility for such systems to host Floquet time crystal phases [43, 44, 45].
+Theoretical insights affirm that symmetry-protected topological order can persist under LR interactions [46] in specific systems.
+However, the generalizability of these findings to broader symmetry-protected topological phases [45, 46, 47, 48, 49, 50] with LR interactions remains an open question, necessitating further investigation.
+These systems give the possibility to study a rich variety of phenomena which contrast the non-equilibrium dynamics of systems with long-range interactions to that of short-range ones, such as the violation of Lieb-Robinson bounds and anomalous propagation of information [51, 52, 53, 54, 55].
+In this paper, we contribute to this burgeoning field by presenting an in-depth study of the quantum Otto engine based on the Ising model (see Fig.2), focusing particularly on the potential influence of quantum criticality on the engine’s thermodynamic performance.Our analysis also reveals that the quantum Otto engine based on the Ising model can function either as a heat engine—extracting work by absorbing heat from the cold reservoir—or as a refrigerator—using input work to transfer heat from the cold to the hot reservoir—depending on the specific parameters governing the thermodynamic cycle.
+While the operational regime is indeed sensitive to these parameters, one can always identify a suitable parameter range in which each mode of operation is achieved with high reliability.
+This study rigorously explores a quantum thermodynamic system governed by the Otto cycle, highlighting its quantum features distinct from classical counterparts [65].
+We first outline the system and cycle, examining the impact of parameters such as system size $N$, initial field $h_{i}$, interaction exponent $\alpha$, and inverse temperatures $\beta_{c}$, $\beta_{h}$, on operational modes, including transitions between refrigerator and engine regimes.
+We then assess the quantum refrigerator and thermal engine, analyzing how $N$ and thermal correlations influence refrigeration capacity $Q_{c}/N$, work output $W/N$, and efficiencies $\eta_{R}$, $\eta$.
+With fixed adiabatic compression time, we investigate idealized machines with infinitely slow phases, quantifying work per spin, efficiency, and stability via the scaling factor $\Pi_{R}/N$ or $\Pi/N$ [65].
+Additionally, we probe correlations for $\alpha\in[0.01,1.3]$ at $N=100$, evaluating efficiency variations with $Q_{c}/N$, $W/N$, and $h_{i}$, and their distribution across chain sizes [30, 31].
+Lastly, we examine the scaling factor proportionality for $N\in[10,100]$, elucidating critical behaviors near phase transitions.
+Figure 1: Dynamical phase diagram for the long-range Ising model ($J=1$), obtained via CMFT with $\ell=4$, following a quench of the transverse field from $h_{\text{in}}=0$ to $h$.
+The plot shows the time-averaged longitudinal magnetization (up to a time $T=1000$) as a function of $\alpha$ and $h$.
+Note that, as soon as $\alpha>1$ (indicated by the dashed line), the dynamical phase transition evolves into a critical region, where the asymptotic magnetization becomes highly sensitive to the system parameters.
+The blue square highlights.
 II quantum otto cycle
@@
 For $\alpha=\infty$, this model reduces to the standard transverse field Ising model (TFIM), which can be exactly solved using the Jordan-Wigner transformation [67, 66] .
-The long range exhibits a quantum phase transition from the ferromagnetic to the paramagnetic phase at $h=2J$.
+The long range exhibits a quantum phase transition from the ferromagnetic to the paramagnetic phase at $h_{c}=J$.
 This transition persists as $\alpha$ decreases [53], with the transition point shifting towards higher magnetic field values.
-At the opposite extreme, $\alpha=0$ [38, 39]represents a fully connected regime, which can be analytically treated for its static and dynamic properties.
-In this case, the model shows long-range ferromagnetic order at low finite temperatures.
-This configuration is particularly interesting for $\alpha<2$, revealing exotic phenomena such as prethermalization [21, 40], dynamical phase transitions [25, 24], and dynamical confinement [41, 42].
-This model is strongly influenced by the parameter $\alpha$.
-For instance, for $\alpha<2$, long-range ferromagnetic order is observed at finite temperatures [43], along with unique dynamic behaviors that have garnered significant experimental interest.
-Trapped ion systems, with their adjustable transverse fields and long-range interactions, provide an ideal framework for studying these phenomena.
-The critical magnetic field is estimated using the Cluster approach and Mean-Field Theory (CMFT) [44, 45, 46] to explore the phase diagram of an Ising spin chain with long-range interactions.
-This method captures short-range correlations while treating interactions between clusters at a mean-field level, revealing the emergence of a chaotic region for specific values of the parameter $\alpha$.
-In our study, we avoid mean-field or cluster decomposition approximations and spin interaction is uniform.
-Nevertheless, we observe significant concordance with the results obtained via CMFT, especially concerning static phase transitions.
-This agreement underscores the robustness of our method, providing an accurate description of phase transitions without resorting to simplifications.
-From this observation, it is possible to extract a qualitative behavior of the quantum critical point as a function of $\alpha$.
-Interpolating the data, we find
-$h_{c}(\alpha)\sim\begin{cases}1&\text{if }\alpha\leq 1,\\ 0.35(3.2-\alpha)&\text{if }1<\alpha<2.\end{cases}$
-Since our method is essentially a generalized mean field theory, we will use it when the latter describes the equilibrium transition ($\alpha<5/3$).
-Despite this, it must be stressed that CMFT is exact for large $\ell$.
-We will use a numerical diagonalization of the Hamiltonian in the disordered case similar to [10], where we will assume $J_{ij}=\frac{1}{K(\alpha)}\frac{J}{|i-j|}$, to compute $\omega_{k}(h)$ for more details consult Otto [35].
-In conclusion, the transverse field Ising model with power-law decaying interactions offers a rich theoretical framework for exploring quantum phase transitions and complex dynamics while being directly applicable to concrete experimental setups.
-II.2 Cycle description
-It describes a one-dimensional long-range Ising system of $N$ quantum spins interacting with a ferromagnetic coupling strength $J>0$, in the presence of magnetic field $h$.
-Hereafter, we set $J=0,5$ as the energy scale and work in units where $\hbar=k_{B}=1$.
+Thus
+$H=-\sum_{j=1}^{N}\sum_{r=1}^{N/2-1}\left[t_{r}c^{\dagger}_{j+r}c_{j}+\Delta_{r}c^{\dagger}_{j+r}c^{\dagger}_{j}+\text{h.c.}\right]-h\sum_{j=1}^{N}\left[1-2c^{\dagger}_{j}c_{j}\right],$ (3)
+where $c^{\dagger}_{j}$ and $c_{j}$ are creation and annihilation operators for fermions at site $j$, while $t_{r}$ and $\Delta_{r}$ are the hopping and pairing amplitudes, respectively.
+We choose their dependence on the intersite distance $r$ according to the power laws
+$t_{r}=\frac{1}{N_{\alpha_{1}}}\frac{J}{r^{\alpha_{1}}},\quad\Delta_{r}=\frac{1}{N_{\alpha_{2}}}\frac{J}{r^{\alpha_{2}}},$ (4)
 Figure 2: The schematic of the Ising Long-range quantum Otto engine consists of two adiabatic phases and two thermalization phases.
@@
 Then, from (C → D), it is adiabatically brought back to $h_{i}$, followed by thermalization at temperature $T_{c}$ in (D → A), completing the cycle.
-The long-range interaction between spins is dictated by the strength parameter $\alpha$ (we set $h_{i}<h_{f}$ and $h_{f}-h_{i}=0.5$).
+with the hopping exponent $\alpha_{1}>0$, the pairing exponent $\alpha_{2}>0$, and $N_{\alpha}=\sum_{r=1}^{N/2}r^{-\alpha}$ the Kac scaling factor [31, 30], which guarantees extensivity of the energy in the case $\alpha_{i}<1$, with $i=1,2$.
+Hereafter, we set $J=\Delta=1$ as the energy scale and work in units of $\hbar=k_{B}=1$.
+This model, and with same range for pairing and hopping $\alpha_{1}=\alpha_{2}=\alpha$ [9].
+Due to the translational invariant nature of the couplings, it is useful to write the Hamiltonian in terms of the momentum-space operators
+$\tilde{c}_{k}=e^{-i\frac{\pi}{4}}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{ikj}c_{j},$ (5)
+where $k=2\pi n/N$, with $n=-N/2+1,\dots,N/2$ (in the following we will drop the $\tilde{}$ on the $c_{k}$ unless it is ambiguous). Then we obtain
+$H=\sum_{k}\left[(h-t_{k})\left(c^{\dagger}_{k}c_{k}-c_{-k}c^{\dagger}_{-k}\right)+\Delta_{k}\left(c^{\dagger}_{k}c^{\dagger}_{-k}+c_{-k}c_{k}\right)\right],$ (6)
+where $t_{k}$ and $\Delta_{k}$ are the Fourier transforms of the hopping and pairing amplitudes, respectively, which in the thermodynamic limit may be written as
+$t_{k}$ $=\frac{\mathrm{Re}\left[\text{Li}_{\alpha_{1}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{1})},$ (7)
+$\Delta_{k}$ $=\frac{\mathrm{Im}\left[\text{Li}_{\alpha_{1}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{1})},$ (8)
+where $\text{Li}_{\alpha}(z)$ denotes the polylogarithm and $\zeta(\alpha)$ is the Riemann zeta function.
+We notice that the Hamiltonian in the Fourier space can be decomposed into the sum of single mode Hamiltonians, introducing $\Psi_{k}=(c_{k},c^{\dagger}_{-k})^{T}$
+$H=\sum_{k}\Psi^{\dagger}_{k}H_{k}\Psi_{k},$ (9)
+$H_{k}=(h-t_{k})\sigma^{z}_{k}+\Delta_{k}\sigma^{x}_{k},$ (10)
+[…]
+$u_{k}=\cos\frac{\theta_{k}}{2},\quad v_{k}=\sin\frac{\theta_{k}}{2},$ (11)
+where $\theta_{k}=\arctan\left[\Delta_{k}/(h-t_{k})\right]$, to obtain
+$H=\sum_{k}\omega_{k}(h)\left(\gamma^{\dagger}_{k}\gamma_{k}-\frac{1}{2}\right),$ (12)
+with the spectrum
+$\omega_{k}(h)=2\sqrt{(h-t_{k})^{2}+\Delta_{k}^{2}}.$ (13)
+At the opposite extreme, $\alpha=0$ , represents a fully connected regime, which can be analytically treated for its static and dynamic properties.
+In this case, the model shows long-range ferromagnetic order at low finite temperatures.
+This configuration is particularly interesting for $\alpha<2$, revealing exotic phenomena such as prethermalization [26, 27], dynamical phase transitions [31, 30], and dynamical confinement [32, 33].
+This model is strongly influenced by the parameter $\alpha$.
+For instance, for $\alpha<2$, long-range ferromagnetic order is observed at finite temperatures [70].
+The critical magnetic field is estimated using the Cluster approach and Mean-Field Theory (CMFT) [71, 72, 73] to explore the phase diagram of an Ising spin chain with long-range interactions.
+This method captures short-range correlations while treating interactions between clusters at a mean-field level, revealing the emergence of a chaotic region for specific values of the parameter $\alpha$.
+In our study, we avoid mean-field or cluster decomposition approximations and spin interaction is uniform.
+Nevertheless, we observe significant concordance with the results obtained via CMFT, especially concerning static phase transitions, it is possible to extract a qualitative behavior of the quantum critical point as a function of $\alpha$.
+Interpolating the data, we find
+$h_{c}(\alpha)\sim\begin{cases}1&\text{if }\alpha\leq 1,\\ 0.35(3.2-\alpha)&\text{if }1<\alpha<2.\end{cases}$
+The nonequilibrium dynamics following a quench may be probed through the time-averaged longitudinal magnetization, defined as $m_{z}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\langle\sigma_{z}(t)\rangle\,dt$, serving as the order parameter.
+Initiating from a fully polarized state, the system undergoes a dynamical quantum phase transition at a critical field $h_{c}=J$.
+In the thermodynamic limit Fig.1, for $0<\alpha<1$, the long-range Ising chain exhibits behavior analogous to the Lipkin-Meshkov-Glick model.
+However, as the power-law exponent surpasses the critical threshold $\alpha=1$, fluctuations become increasingly pivotal, profoundly shaping the collective dynamics of the system.
+Since our method is essentially a generalized mean field theory, we will use it when the latter describes the equilibrium transition ($\alpha<5/3$).
+Despite this, it must be stressed that CMFT is exact for large $\ell$.
+We will use a numerical diagonalization of the Hamiltonian in the disordered case (Consult appendix  A and Bfor more details) .similar to [15], where we will assume $J_{ij}=\frac{1}{K(\alpha)}\frac{J}{|i-j|}$, to compute $\omega_{k}(h)$ for more details consult Otto [66].
+In conclusion, the transverse field Ising model with power-law decaying interactions offers a rich theoretical framework for exploring quantum phase transitions and complex dynamics while being directly applicable to concrete experimental setups.
+II.2 Cycle description
 We will work on an ideal otto cycle illustrated in Fig.2 composed of two infinitely slow adiabatic phases and two thermalization phases without any dissipation.
@@
 For the dynamic and ideal case, we have:
-A → B: Adiabatic increase of the transverse field
-The field $h(t)$ increases linearly from $h_{i}$ to $h_{f}$, while the working substance is kept isolated from the thermal baths. The expression for $h(t)$ is:
-$h(t)=h_{i}+0.5\frac{t}{T}$
-  • For the ideal case: $T\to\infty$ and $h=h_{i}$
-B → C: Thermalization with the cold bath
-Once the field reaches $h_{f}$, the magnetic field is fixed, and the working substance is brought into contact with the cold bath until it reaches thermal equilibrium at temperature $T_{c}$ (we will assume a perfect thermalization).
-C → D: Adiabatic decrease of the transverse field
-The field $h(t)$ is decreased linearly from $h_{f}$ back to $h_{i}$ at the same rate as in step A → B, while keeping the working substance isolated from the thermal baths.
-The expression for $h(t)$ is:
-$h(t)=h_{f}-0.5\frac{t}{T}$ (3)
-  • For the ideal case: $T\to\infty$ and $h=h_{f}$
-D → A: Thermalization with the hot bath.
-Once the field returns to $h_{i}$ is kepet fixed, and the working substance is brought into contact with the hot bath until it returns to its initial thermal state at temperature $T_{h}$.
+  • (a)
+    $A\to B$: Adiabatic field increase;
+    The transverse field is linearly ramped up from $h_{i}$ to $h_{f}$ over time, while the working substance remains isolated from the thermal baths.
+    $h(t)=h_{i}+vt,\quad t\in\left[0,\frac{h_{f}-h_{i}}{v}\right]$
+  • (b)
+    $B\to C$: Thermalization with the cold bath;
+    The Hamiltonian $\hat{H}_{\text{sys}}(t_{f})$ is held constant, and the system is coupled to the cold bath until the working substance reaches the thermal state of $\hat{H}_{\text{sys}}(t_{f})$ at temperature $T_{c}$.
+  • (c)
+    $C\to D$: Adiabatic field decrease;
+    The transverse field is linearly reduced from $h_{f}$ back to $h_{i}$ at the same rate as in step (a), with the working substance decoupled from the baths.
+    $h(t)=h_{f}-vt,\quad t\in\left[0,\frac{h_{f}-h_{i}}{v}\right]$
+  • (d)
+    $D\to A$: Thermalization with the hot bath;
+    The Hamiltonian $\hat{H}_{\text{sys}}(t_{i})$ is maintained, and the system is connected to the hot bath until the working substance returns to its initial thermal state of $\hat{H}_{\text{sys}}(t_{i})$ at temperature $T_{h}$.
+Figure 3: The color code represents the following modes: green for accelerator, purple for heater, red for heat engine, and blue for refrigerator.
+quantities based on the Appendix A [74, 75]. The efficiency of the cycle can thus be expressed as:
+$\eta=\frac{W}{Q_{h}}.$ (14)
+In what follows, we use $Q_{c}(h)$ to represent the heat exchanged during thermalization with the cold (or hot) bath, adopting the convention that $Q_{c}(h)>0$ when the system absorbs heat from the reservoir.
+If $\rho_{\alpha}$ is the density matrix of the system at points $\alpha=A,B,C,D$, then:
+$Q_{c}$ $=\langle\hat{H}(t_{f})\rangle_{\rho_{C}}-\langle\hat{H}(t_{f})\rangle_{\rho_{B}},$ (15)
+$Q_{h}$ $=\langle\hat{H}(t_{i})\rangle_{\rho_{A}}-\langle\hat{H}(t_{i})\rangle_{\rho_{D}}.$ (16)
+It describes a one-dimensional long-range Ising system of $N$ quantum spins interacting with a ferromagnetic coupling strength $J>0$, in the presence of magnetic field $h$.
+Hereafter, we set $J=1.0$ as the energy scale and work in units where $\hbar=k_{B}=1$.The long-range interaction between spins is dictated by the strength parameter $\alpha$ (we set $h_{i}<h_{f}$ and $h_{f}-h_{i}=0.5$).
 When analyzing the case of an infinitely slow cycle, i.e., the limit $T\to\infty$, the regime is typically referred to as adiabatic.
 In this regime, the unitary evolution is slow enough to satisfy the adiabatic theorem, preventing transitions between the instantaneous eigenstates of the Hamiltonian.
-We will calculate the following quantities based on the Appendix A [47, 48, 49].
-The efficiency of the cycle can thus be expressed as:
-$\eta=\frac{W}{Q_{h}}.$ (4)
-For an ideal Otto cycle, strictly speaking, perfect thermalization and perfect quantum adiabatic sweeps are ideally obtained in the limits $T\to\infty$, can be written as follows:
+The nonlocal master equation chosen to model the system–environment interaction [56].
+Let us consider, for a while, the more general configuration of $N_{B}$ independent thermal reservoirs at temperature $T_{n}$, with $n\in\{1,\dots,N_{B}\}$ indices labeling the bath.
+The Hamiltonian describing this setup reads
+$\hat{H}_{\text{env}}=\sum_{n=1}^{N_{B}}\int dk\,\varepsilon_{n}(k)\hat{c}_{n}^{\dagger}(k)\hat{c}_{n}(k),$
+with $\varepsilon_{n}(k)\geq 0$ and $\hat{c}_{n}(k)$, $\hat{c}_{n}^{\dagger}(k)$ being fermionic annihilation and creation operators.
+The $N_{B}$ baths are independent, therefore the corresponding reduced density operator of the full environment assumes the factorized form
+$\rho_{\text{env}}=\bigotimes_{n=1}^{N_{B}}\rho_{\text{bath}}^{(n)},$
+where $\rho_{\text{bath}}^{(n)}$ is the thermal density matrix describing the $n$th fermionic bath at temperature $T_{n}$.
+[…]
+Tracing out all the environmental degrees of freedom and imposing the Born–Markov approximation for the baths, it is possible to derive a microscopic Lindblad master equation , in the energy eigenbasis, for the reduced density matrix of the system described by the Hamiltonian :
+$\frac{d}{dt}\rho_{\text{sys}}(t)=-i[\hat{H}_{\text{sys}},\rho_{\text{sys}}]+D[\rho_{\text{sys}}],$
+$D[\rho_{\text{sys}}]=\sum_{n,k}\gamma_{nk}$ $\bigg{[}(1-f_{n}(\omega_{k}))\left(\frac{1}{2}\hat{b}_{k}\rho_{\text{sys}}\hat{b}_{k}^{\dagger}-\frac{1}{2}\{\hat{b}_{k}^{\dagger}\hat{b}_{k},\rho_{\text{sys}}\}\right)$ (17)
+$+f_{n}(\omega_{k})\left(\frac{1}{2}\hat{b}_{k}^{\dagger}\rho_{\text{sys}}\hat{b}_{k}-\frac{1}{2}\{\hat{b}_{k}\hat{b}_{k}^{\dagger},\rho_{\text{sys}}\}\right)\bigg{]}.$
+where the $\{\hat{b}_{k},\hat{b}_{k}^{\dagger}\}$ jump operators are the fermionic Bogoliubov quasiparticles which diagonalize the model , these operators are local in the energy eigenbasis and thus nonlocal in the sites, giving rise to a global master equation [56].
+Moreover,
+$f_{n}(\omega_{k})=\frac{1}{1+e^{\omega_{k}/T_{n}}},$
+Under the assumption of no degeneracies in the spectrum (as turns out to be the case for the Ising chain with open boundary conditions), it can be used to obtain an analytic expression for the time evolution of the correlation functions.
+In particular, defining
+$\tilde{f}_{k}=\frac{\sum_{n}\gamma_{nk}f_{n}(\omega_{k})}{\sum_{n}\gamma_{nk}},$
+we have
+$\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle_{t}=\tilde{f}_{k}\left(1-e^{-2\sum_{n}\gamma_{nk}t}\right)+\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle_{0}e^{-2\sum_{n}\gamma_{nk}t}.$
+Then, all the thermodynamic quantities can be easily computed using equation (20) or, equivalently, directly computing the partition function $Z$. In particular, the internal energy reads:
+$\langle E_{i}\rangle=\sum_{k}\omega_{k,i}\left(\langle b_{k}^{\dagger}b_{k}\rangle-\frac{1}{2}\right).$ (18)
+For an ideal Otto cycle, in order to avoid any complications arising from the irreversibility inherent to quantum friction , we shall consider perfectly adiabatic strokes by assuming an infinitely slow transformation time $t\to\infty$ , This regime is usually referred to as adiabatic, since the unitary evolution is sufficiently slow for the adiabatic theorem to hold, preventing transitions between the instantaneous eigenstates of the Hamiltonian , which leads us back to the Fermi–Dirac distribution;
+$\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle_{t\to\infty}=\tilde{f}_{k}.$
+The thermodynamic quantities are calculated as follows:
 $Q_{h}=\sum_{k}\omega_{k}(h_{f})\Delta f_{k},$
@@
 $\eta=\frac{W}{Q_{h}}.$ (19)
+where we set
+$\Delta f_{k}\equiv f(\beta_{h},\omega_{k}(h_{f}))-f(\beta_{c},\omega_{k}(h_{i})).$ (20)
 To calculate the efficiency $\eta_{R}$ of the refrigerator, defined as
 $\eta_{R}=\frac{Q_{c}}{|W|}.$ (21)
-where
-$\Delta f_{k}\equiv f(\beta_{h},\omega_{k}(h_{f}))-f(\beta_{c},\omega_{k}(h_{i})).$ (7)
+Figure 5: The influence of long-range interactions on the work and the quantum refrigerator’s performance per spin as a function of $h_{i}$.
+Here NN is fixed at 100, while $\alpha$ is varied from 0 to $1.3$.
+Figure 6: In the figure, NN is fixed at 100, and $\alpha$ is varied from 0 to 1.3 to observe the role of long-range interactions on the efficiency $\eta$ of the quantum heat engine and the coefficient of performance $\eta_{R}$ of the refrigerator as a function of $h_{i}$.
 Note that, since after a single ideal cycle the system returns to the same initial state, we have $\Delta E=0$ and thus $W=Q_{h}+Q_{c}$ follows from the first law of thermodynamics.
 II.3 Operations modes
-The nature of our quantum heat engine is dictated by the sign of the following quantities $Q_{c}$, $Q_{h}$, and $W$.
-As illustrated in Fig.(2), our engine can operate in one of the following four modes:
-  • E
-    (Heat engine): $Q_{c}\geq 0,Q_{h}\leq 0,W\geq 0$, the engine produces work by absorbing heat from the hot reservoir.
-  • R
-    ( Refrigerator): $Q_{c}\leq 0,Q_{h}\geq 0,W\leq 0$, the engine removes heat from the cold reservoir and expels it to the hot reservoir by converting the absorbed work.
-  • A
-    (Accelerator): $Q_{c}\geq 0,Q_{h}\leq 0,W\leq 0$, the engine absorbs energy and heats the cold reservoir from the hot reservoir.
-  • H
-    (Heater): $Q_{c}\leq 0,Q_{h}\leq 0,W\leq 0$, the engine absorbs heat and heats both reservoirs.
-Figure 2: The color code represents the following modes: green for accelerator, purple for heater, red for heat engine, and blue for refrigerator.
-Figure 3: The different operating regimes of the quantum Otto engine in the $\alpha,\beta_{h}/\beta_{c},N-h_{i}$ plane.
-The color code represents the following modes: accelerator (green), heater (purple), heat engine (red), and refrigerator (blue).
-The purpose of this section is to show the switching between operational modes and their dependencies on the number of particles $N$, the interaction strength $\alpha$ , and the ratio $\beta_{h}/\beta_{c}$.
+The purpose of this section is to show the switching between operational modes and their dependencies on the number of particles $N$, the interaction strength $\alpha$ , and the ratio $\beta_{h}/\beta_{c}$.Let us now analyze the case of an infinitely slow cycle , the limit $t\to\infty$.
 All $\alpha$ values can be considered as long-range correlations, but for comparison purposes, we will divide them into three categories: long-range, intermediate, and short-range.
-This is because, as we increase $\alpha$, the effect of long-range correlations decreases and weakens.
-We have chosen specific values of $\alpha$, $\beta_{h}$, and $\beta_{c}$ where a diversity of modes can be observed in the ideal case.
-Considering the specific values imposed on our cycle, we observe the following behaviors:
-For Fig.3($a$), where the system size is fixed at $N=100$, with $\beta_{h}=1.76$, $\beta_{c}=2.65$, and where $\alpha$ varies from 0 to 0.6, we observe different operational modes of the quantum engine as a function of the initial magnetic field $h_{i}$.
-For low values of $\alpha$ (between 0 and 0.2), long-range correlations between particles impose a refrigerator mode (blue) before the critical field $h=1$.
-For intermediate fields ($h_{i}=1.2$ to $1.7$), the accelerator (green) and heater (purple) modes appear; we see that they shift near the critical point to higher fields, playing the role of transient modes for the engine mode (red), which is predominant at high fields ($h_{i}=1.5$ to $2.5$).
-For intermediate $\alpha$ values (between $0.2$ and $0.4$), we observe a significant decrease in the refrigerator mode, indicating that medium-range correlations favor the appearance of the engine mode on both sides of the critical point ($h_{i}=0.5$ to $1.5$).
-The transient modes shift from $h_{i}=1.2$ at $\alpha=0.2$ to $h_{i}=0.2$ at $\alpha=0.4$.
-At intermediate fields, the heater and accelerator modes dominate, extending to high fields, with a significant decrease in the engine mode and the emergence of a refrigerator mode at high fields.
-For short-range correlations ($\alpha$ between $0.4$ and $0.6$), there is a slight decrease in the refrigerator mode and the disappearance of the transient modes A (accelerator) and H (heater), while the engine mode increases as $\alpha$ goes from $0.4$ to $0.6$.
-The H, A, and R (refrigerator) modes shift toward higher fields.
-The figure illustrates the impact of the magnetic field and correlation range on the operational modes.
-In Fig.3($b$) the operational behavior is studied as a function of the ratio $\beta_{h}/\beta_{c}$, ranging from $0.45$ to $0.75$, with $\beta_{c}$ varying between 2.345 and 3.911, for a system size of $N=100$, with $\beta_{h}=1.76$ and $\alpha=0.25$.
-For higher values of $\beta_{h}/\beta_{c}$ (between $0.65$ and $0.75$), the refrigerator mode (blue) dominates and extends around the critical field at $h_{i}\approx 1.0$.
-The accelerator (A) and heater (H) modes also appear, acting as transitional modes between lower fields and the heat engine (E) in regions of more intense fields.
-As the $\beta_{h}/\beta_{c}$ ratio decreases (between 0.65 and 0.55), the refrigerator mode shows a significant reduction, while the A and H modes expand around the critical field, taking over from the high-field regions.
-Meanwhile, the zone corresponding to the engine mode (E) grows in the high-field regions and, to a lesser extent, in pre-transition phase chains.
-When the $\beta_{h}/\beta_{c}$ ratio is particularly low (between 0.45 and 0.55), the refrigerator mode rapidly declines for low fields, while the heater mode almost completely disappears.
-The accelerator mode, on the other hand, emerges after the phase transition, around $h_{i}\approx 1.5$, while the engine mode (E) strengthens, covering both low and high fields.
-These observations indicate that as the $\beta_{h}/\beta_{c}$ ratio approaches 1, the system tends to operate in refrigerator mode.
-However, as this ratio decreases, there is a gradual transition to heat engine mode, marked by the appearance and reinforcement of the accelerator and heater modes.
-Fig.3($c$) examines the operational zones as a function of the system size $N$, which varies from $10$ to $100$, with $\alpha=0.25$, $\beta_{h}=1.76$, and $\beta_{c}=2.65$.
-For small system sizes (from $10$ to $40$), the refrigerator mode (blue) shows a dominance of the engine mode (red) for low fields before the critical point, with the appearance of the refrigerator mode at $N=20$ for low $h_{i}$.
-For intermediate fields, the accelerator (green) and heater (purple) modes appear (from $h_{i}=1.5$ to $2$), and the system operates exclusively in the refrigerator mode for strong fields.
-As we move to medium-length chains (from $40$ to $70$), there is a slight increase in the refrigerator mode for low $h_{i}$, with the appearance of the accelerator and heater modes and a slight retreat of the engine mode around the critical field.
-We also observe a continuous migration of the accelerator and heater modes from pre-critical fields toward the critical field, favoring the refrigerator mode, which appears in the intermediate and strong fields.
-For long chains (from $70$ to $100$), there is an increase in the refrigerator mode, which extends from low $h_{i}$ up to the critical field, with the disappearance of the engine mode after $N=92$ for low to intermediate fields.
-Additionally, the heater and accelerator modes expand from intermediate to strong fields, while the refrigerator mode barely appears and disappears in strong magnetic fields, favoring the engine, heater, and accelerator modes.
-We observe that fixing the range of interactions and the reservoir temperatures does not stabilize the operational modes since they also depend on the chain size.
-This analysis reveals the intricate interplay between system size, interaction strength, and temperature ratio $\beta_{h}/\beta_{c}$ in determining the operational modes of a quantum engine.
-The results demonstrate that the operational modes are influenced not only by the chosen values for $\alpha$, $\beta_{h}$, and $\beta_{c}$, but also by the system size $N$.
-Through the analysis of the figures, it is evident that long-range correlations favor the refrigerator mode at low values of $\alpha$, while at intermediate values, the accelerator and heater modes become predominant.
-Additionally, the transient modes vary depending on the applied magnetic fields, illustrating complexdynamics when crossing the critical point.
-It has also been shown that varying the ratio $\beta_{h}/\beta_{c}$ significantly impacts the operational modes, favoring the refrigerator mode at higher values and leading to a transition to the heat engine mode at lower values.
-These results underscore the importance of a thorough understanding of interactions and correlations within the system to optimize the performance of quantum engines.
-This study paves the way for further explorations in optimization of the parameters to effectively exploit the different operational modes for practical applications in the field of quantum energy.
+The color code represents the following modes: accelerator (green), heater (purple), heat engine (red), and refrigerator (blue).As illustrated in Fig.3, our engine can operate in one of the following four modes:
+  • Refrigerator (R): the engine absorbs energy and transfers heat from the cold reservoir to the hot one, i.e., $Q_{c}>0$, $Q_{h}<0$, and $W<0$;
+  • Accelerator (A): the engine absorbs energy and transfers heat from the hot reservoir to the cold one, i.e., $Q_{c}<0$, $Q_{h}>0$, and $W<0$;
+  • Heat engine (E): the engine produces work by absorbing heat from the hot reservoir, i.e., $Q_{c}<0$, $Q_{h}>0$, and $W>0$;
+  • Heater (H): the engine absorbs energy and heats up both the hot and the cold reservoirs, i.e., $Q_{c}<0$, $Q_{h}<0$,and $W<0$.
+Figure 4-a summarises the machine’s operational regimes for $N=100$, $T_{h}=0.57$ and $T_{c}=0.38$ as the interaction range is varied.
+Physically, decreasing $\alpha$ (making the interactions longer-ranged) strengthens collective correlations and coherent energy exchange across the chain: this favors refrigeration at low fields and promotes nonlocal response modes at intermediate fields, because the system can coordinate energy flows across many sites.
+By contrast, increasing $\alpha$ (shortening the interaction range) suppresses these collective channels, which reduces the region where refrigeration and transient modes appear and instead stabilises steady engine-like operation over a wider field window.
+In other words, longer-range couplings enhance sensitivity to global field conditions and produce pronounced finite-$N$ modes, whereas more local couplings favour robust, local energy extraction and smoother, more extensive engine performance.
+The former case $\alpha=0.2$, lies well outside the perturbative domain of Fisher’s expansion and corresponds to an interaction range so extended that non-additive, quasi–mean-field behaviour is expected; the latter, $\alpha=1.2$, remains within the classical long-range sector but with markedly weaker long-range effects.
+These conclusions are therefore fully compatible with the analytical expectations of [10] and with the qualitative differences observed in our numerical data.
+Figure 4-b displays the machine’s operational map for $N=100$ and $\alpha=0.25$, with $T_{h}=0.57$ while the cold temperature is varied so that the ratio $T_{c}/T_{h}$ spans 0.450.450.750.75.
+Physically, changing $T_{c}/T_{h}$ tunes the thermodynamic bias available for work extraction: larger ratios (temperatures closer together) reduce the net thermal gradient and suppress steady work output, making thermally-driven refrigeration and transient response modes more prominent near the critical field.
+Conversely, decreasing $T_{c}/T_{h}$ (increasing the thermal contrast) strengthens the directional heat flow that drives net work production, stabilising engine-like operation across a broader field range and diminishing the parameter window where refrigerator or heater modes appear.
+Transient accelerator/heater bands that flank the critical region reflect a competition between critical susceptibility (which amplifies small driving imbalances) and the global thermal bias; as the thermal bias grows, this competition is resolved in favour of persistent engine behaviour.
+Figure 4-c shows how the machines operating regimes reorganize as the system size $N$ is varied (here $\alpha=0.25$, $T_{h}=0.57$, $T_{c}=0.38$).
+Physically, the observed changes reflect a competition between two size-dependent effects: (i) finite-$N$ coherence and level discreteness, which favour directed, coherent energy extraction when the chain is short, and (ii) the progressive densification of the many-body spectrum and the growth of low-energy collective modes as $N$ increases, which amplify critical susceptibility and thermal redistribution near the critical field.
+For small chains, the sparsity of excitations and enhanced coherence make steady engine-like operation energetically favourable at low fields.
+As $N$ grows the spectrum becomes denser and low-energy fluctuations proliferate; these modes couple more efficiently to the reservoirs and to field-induced critical responses, expanding refrigeration and transient accelerator/heater bands close to criticality and reducing the parameter window for steady engine operation.
+At the largest sizes the increased phase-space for scattering and thermal equilibration further reinforces refrigeration-dominated behaviour over a broad field interval.
+In short, increasing $N$ shifts the device from a regime dominated by finite-size coherent extraction to one governed by many-body critical fluctuations and thermal redistribution, which explains the progressive replacement of engine by refrigerator behaviour in the data.The device’s operational mode is set by three parameters: the interaction exponent $\alpha$, the thermal contrast $T_{c}/T_{h}$ (or $\Delta T$), and the system size $N$.
+Smaller $\alpha$ (long-range couplings) enhances global coherence and favours refrigeration with pronounced finite-$N$ transients near criticality, whereas larger $\alpha$ suppresses collective channels and stabilises engine-like operation.
+A small thermal gap (high $T_{c}/T_{h}$) weakens net work extraction and biases the system toward refrigeration or transient responses, while a large gap strengthens sustained engine performance.
+Small $N$ permits coherent finite-size windows for directed extraction; large $N$ increases low-energy fluctuations and thermalisation phase-space, broadening refrigeration and transient regimes.
+Hence $\alpha$, $T_{c}/T_{h}$, and $N$ constitute direct, tunable levers for steering and optimising quantum thermal machines.
+Figure 7: Efficient $\eta$ (COP $\eta_{R}$ ) as a function of $h_{i}$ and W/N ($Q_{c}$/N) this highlights the behavior of the efficiency for different initial field values $h_{i}$ and of the work per spin $W/N$, as well as the possible optimization paths of the efficiency for various interaction ranges characterized by different values of $\alpha$.
+.
 II.4 Long Range Impact
-Fig.4($a$) presents an analysis of the operational modes of a quantum thermal engine and a refrigerator in the ideal case.
-Initially, with $N$ fixed at $100$ and $\alpha$ varied from $0.01$ to $1.3$, we investigate the influence of interaction range and correlations on the efficiency and work output per particle $W/N$ of the heat engine, and on the efficacy and extracted heat per particle $Q_{c}/N$ of the refrigerator.
-Subsequently, we analyze the efficiency ($\eta$, $\eta_{R}$) as a function of $W/N$ and $Q_{c}/N$, respectively, for both the heat engine and refrigerator, considering system sizes $N$ ranging from $10$ to $100$.
-We analyze the different operational modes of a quantum heat engine and refrigerator in the ideal case; For the refrigerator, we consider $\beta_{h}=1.4$ and $\beta_{c}=1.5$ ($\beta_{h}=0.58$ and $\beta_{c}=0.6$), while for the thermal machines, $\beta_{h}=0.2$ and $\beta_{c}=0.9$ ($\beta_{h}=0.2$ and $\beta_{c}=0.667$).
-Figure 4: The influence of long-range interactions on the work and the quantum refrigerator’s performance per spin as a function of $h_{i}$.
-Here $N$ is fixed at 100, while $\alpha$ is varied from 00 to $1.3$.
-Figure 5: In the figure, $N$ is fixed at 100, and $\alpha$ is varied from 0 to 1.3 to observe the role of long-range interactions on the efficiency $\eta$ of the quantum heat engine and the coefficient of performance $\eta_{R}$ of the refrigerator as a function of $h_{i}$.
-The ideal thermal machine is presented in the two figures, Fig.4($a$-$b$) and Fig.5($a$-$b$), where $W/N$ and $W/Q_{h}$ are again plotted as a function of $h_{i}$ for different ranges of $\alpha$.
-In this idealized scenario, we observe that for reservoirs with a large difference in reservoir temperatures (with $\beta_{h}=0.2$ and $\beta_{c}=0.667$), higher values of $\alpha$, which favor shorter correlations, result in higher work per spin $W/N$ compared to smaller and intermediate $\alpha$ before the critical point ($0.66$ for $\alpha=1.0$ to $0.77$ for $\alpha=1.3$).
-There is a significant increase in $W/N$ for larger $h_{i}$, and near the critical field, a saturation of $W/N$ values occurs for larger $\alpha$, showing similar behavior.
-In the post-critical field region, an inversion occurs, with lower $\alpha$ ranges showing better $W/N$ than intermediate and higher ranges.
-This indicates that post-critical fields tend to favor dynamics that are more advantageous for short-range correlations.
-When the thermal gap is reduced (with $\beta_{h}=0.2$ and $\beta_{c}=0.9$), we observe higher work per spin for long-range correlations at very low magnetic fields, with a transition occurring before the phase transition, for short-range correlations, the work per spin $W/N$ improves, with a peak observed around $h_{i}=1$ at phase transition for long range correlations.
-The efficiency $\eta$ exhibits similar behavior, independent of the thermal gap.
-It is higher for long-range correlations, with a pronounced peak at low $h_{i}$.
-The thermal gap slightly shifts the peak but affects all correlations in the same way (see Fig.5($a$-$b$)).
-Let us now move on to the same study for the refrigerator in the case where the reservoires temperature difference is shorter, with $\beta_{h}=1.4$ and $\beta_{c}=1.5$, a similar pattern is observed for $\frac{Q_{c}}{N}$ , with greater work per spin for long-range correlations.
-The distinction becomes more evident in the efficiency $\eta_{R}$, which shows a marked rise for all $\alpha$ values, especially the higher ones.
-A notable saturation occurs at the critical point, causing a disruption and a decline in short-range correlations at $h_{i}=1.5$, rendering them less effective than long-range ones.
-At the same time, intermediate $\alpha$ values stabilize, experiencing a slight decrease but still outperforming the higher and short $\alpha$ values (red, blue).
-In the case of the ideal refrigerator, Fig.4($c$-$d$) and Fig.5($c$-$d$), the first two figures depict the behavior where the ratio $\eta_{R}=\frac{Q_{c}}{|W|}$ and the coefficient of performance $\frac{Q_{c}}{N}$ are plotted against $h_{i}$ for different values of $\alpha$, with two very close reservoirs ($\beta_{h}=0.58$ and $\beta_{c}=0.6$).
-We observe that for low $h_{i}$ values (before $h_{i}=0.3$), short- and intermediate-range correlations are favored, allowing more heat to be extracted from the cold reservoir per spin compared to smaller $\alpha$.
-For slightly higher magnetic fields (from $h_{i}=0.3$ to $h_{i}=1$), there is a shift, and long-range correlations (small $\alpha$) begin to dominate.
-These long-range correlations saturate particularly near the critical field.
-After the phase transition, a significant increase in $\eta_{R}$ and $\frac{Q_{c}}{N}$ is observed for all $\alpha$ values, without a reversal.
-Starting from $h_{i}=1.5$, stabilization occurs, with a slight decrease in performance for large and intermediate $\alpha$, while small $\alpha$ maintains its effectiveness, indicating that long-range correlations are more favorable in the post-critical region.
-In an ideal refrigerator, we also observe that for small values of $\alpha$ (e.g., from 0.1 to 0.4), both $\frac{Q_{c}}{|W|}$ and $\frac{Q_{c}}{N}$ gradually increase with $h_{i}$, reaching a peak for higher values of $\alpha$.
-As $\alpha$ increases, the system is able to maintain its cooling efficiency over a wider range of $h_{i}$.
-This suggests that, under ideal conditions, the refrigerator can sustain a stable cooling effect even as the range of interactions extends, with improved performance observed for higher values of $\alpha$.
-Figure 6: Efficient $\eta$ (COP $\eta_{R}$) as a function of $h_{i}$ and W/N ($Q_{c}$/N).
-In Fig.6($c$-$d$), which illustrates the effective efficiency $\eta_{R}$ as a function of $Q_{c}/N$ and the initial transverse field $h_{i}$, small values of $N$ (around 10-30) exhibit a high sensitivity to variations in $\frac{Q_{c}}{N}$ and $h_{i}$.
-The effective efficiency $\eta_{R}$ reaches elevated values within a narrow range of $Q_{c}/N$, indicating an unstable behavior of the system for these smaller sizes.
-As $N$ increases, the curves become smoother, suggesting a stabilization of the efficiency and a reduction in sensitivity to fluctuations.
-For higher values of $N$, the maximum value of $\eta_{R}$ shifts towards lower $Q_{c}/N$, indicating that optimal efficiency is achieved with lower heat extraction per cycle, likely due to a more effective management of interactions in a larger system.
-We observe a sequential behavior: before the phase transition, we have higher efficiency $\eta_{R}$ for short chains and larger heat extraction $Q_{c}/N$ from the cold reservoir for long chains.
-At the critical field $h_{i}=1$, a shift occurs, and the behavior changes.
-Long chains become more efficient, with $\eta_{R}$ exceeding that of short chains, and $Q_{c}/N$ becomes significant for small $N$, indicating an $N$-dependence of the efficiency after the phase transition.
-For the second case with $\alpha=1.2$, we observe a distinct behavior and greater stability compared to the previous case.
-This indicates that as $\alpha$ increases, short-range correlations are favored, making them less sensitive to thermal fluctuations.
-Consequently, they become less dependent on the phase transition, which primarily enhances efficiency and $Q_{c}/N$, both of which are more significant for short chains.
-In Fig.7($a$-$b$) In the following, we will comment on the figure, which represents the thermodynamic efficiency $\eta$ plotted as a function of $h_{i}$ and the work per spin $\frac{W}{N}$.
-For small values of $N$ ($10$ to $40$), at low $h_{i}$, we observe a significant efficiency, although lower than that of intermediate and longer chains, with a higher work per spin for smaller $N$.
-As the magnetic field increases, there is a sharp drop for short chains near the critical field $h=1$, with a slight decrease for larger $N$, which increases fluctuations for small $N$ and quickly lowers $\frac{W}{N}$ for stronger chains.
-Meanwhile, for longer and intermediate chains, at the critical field, there is a drop in efficiency and a significant increase in the work per spin $\frac{W}{N}$.
-This indicates that the efficiency of the cycle strongly depends on the initial applied field and the size of the chain.
-In this study, we have explored the operational modes of a quantum heat engine and a refrigerator, emphasizing the effects of long- and short-range correlations as governed by the parameter $\alpha$.
-By examining various values of $\alpha$ and different reservoir temperature configurations, we have identified distinct behaviors in both the quantum thermal machine and the refrigerator.
-For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the pre-critical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region.
-This demonstrates that critical dynamics significantly impact the work per spin $W/N$ and the efficiency.
-In the case of larger reservoir temperature differences, $\alpha$ plays a crucial role in determining performance, especially near the critical point where saturation is observed.
-For the refrigerator, long-range correlations are similarly favored post-critical field, particularly in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$.
-The proximity of the reservoirs’ temperatures reveals that long-range correlations maintain stable performance even after the critical transition, while intermediate and short-range correlations exhibit a decline in efficiency.
-Overall, the analysis highlights the significant influence of correlation range and magnetic field strength $h_{i}$ on the thermodynamic properties of both systems.
-Long-range interactions provide distinct advantages in certain regimes.
-Moreover, the stabilization of these effects for larger system sizes $N$ suggests that larger chains offer better optimization of quantum thermal devices, while smaller chains are more sensitive to fluctuations, particularly near the critical field.
-This study provides valuable insights into optimizing quantum heat engines and refrigerators through the control of system parameters and interactions.
+In the present study, we explore how the distinctive attributes of long-range interacting quantum systems can enhance the efficiency of a many-body quantum Otto cycle, adopted here as a representative model of a quantum thermal device.
+Our analysis reveals several key benefits similar to [9], derived from the incorporation of long-range interactions:
+  1. 1.
+    The infinitely slow limit , the device attains markedly improved optimal performance.
+This enhancement is most evident in the practically relevant heat-engine and refrigerator regimes when collective and local effects are optimally balanced.
+.
+  2. 2.
+    The interaction-range exponent $\alpha$ affects the cycle efficiency differently depending on the operational mode.
+Relying on the results reported by [11], it was shown that the spatial decay of connected correlations depends sensitively on $\alpha$ and on proximity to the critical phase transition: large $\alpha$ yields short-range exponential screening, small $\alpha$ produces algebraic long-range tails, and an intermediate hybrid crossover combines both behaviours.
+On the paramagnetic side, quasi-exponential screening amplifies nearest-neighbour influence, whereas sufficiently small $\alpha$ restores system-wide coherence and collective channels.
+Crucially, the presence of a spectral gap does not preclude algebraic tails, so that spectral structure intrinsically limits the efficacy of nonlocal propagation.
+In Figures 5 and 6, thermodynamic performance is plotted as a function of the initial magnetic field $h_{i}$ for various values of $\alpha$.
+When the temperature gap is large ($T_{h}=5.0$, $T_{c}=1.11$) , an increase in $\alpha$, corresponding to a reduction in the range of correlations, leads to a higher work output per spin before the critical region.
+At the critical phase, an inversion occurs where long-range interactions with smaller $\alpha$ outperform those with larger $\alpha$, suggesting that post-critical dynamics favor long-range interactions.
+This is explained by the influence of neighboring particles, which is more effective in the paramagnetic phase, where increased spin misalignment enhances work extraction.
+When the thermal gap is reduced (e.g., $T_{c}=0.5$, $T_{h}=1.67$), long-range interactions yield a higher $W/N$, as an increase in $\alpha$ induces localized interactions that propagate more weakly than long-range interactions when the thermal gap is small.
+The efficiency $\eta$ exhibits a similar trend across configurations, reaching a maximum for long-range correlations at low $h_{i}$, with a slight shift in the peak depending on the thermal gap.
+Finally, in the refrigerator configuration with a narrow thermal gap ($T_{h}=0.71$, $T_{c}=0.67$ and $T_{h}=1.72$, $T_{c}=1.67$), a similar behavior is observed for the cooling power $Q_{c}/N$, with long-range interactions yielding superior performance, as the extended range of interactions facilitates more effective heat diffusion compared to larger $\alpha$ values, which prioritize localized transmission.
+An increase in $\alpha$ for the efficiency $\eta_{R}$, which reduces dissipation to nearest neighbors, significantly enhances the yield until it reaches a saturation threshold.
+Guided by the conclusions of [12, 13], the observed saturation of the critical peaks is a direct consequence of finite-temperature rounding: the thermal length scale $\xi_{T}$ bounds the spatial correlations, so that finite-size scaling applies only for $N\lesssim\xi_{T}$, whereas for $N\gg\xi_{T}$ would-be critical divergences are suppressed and instead manifest as extended plateaus in the relevant observables.
+Long-range couplings (small $\alpha$) thus enhance performance for short chains but produce rapid saturation once $N$ exceeds $\xi_{T}$; by contrast, larger $\alpha$ yields a more gradual increase in performance that ultimately levels off.
+Namely the coexistence in the QLRO phase of violations of the area law for entanglement entropy and a gapped entanglement spectrum are consistent with a picture in which the spread of quantum correlations is limited.
+This limited spread explains the saturation of the performance peaks we observe and suggests the existence of an energetic barrier that naturally limits the spread of quantum correlations.
+Figure 8: Contour plot of $Q_{c}$ denotes the the heat exchanged during the thermalisation with the cold bath done by the system and $\eta_{R}$ as a function of $h_{i}$ for different values of $\alpha$, with (a) $\alpha=0.25$, (b) $\alpha=1.2$, (c) $\alpha=0.25$, and (d) $\alpha=1.2$.
+Figure 9: COP of the refrigerator versus $h_{i}$, for different system sizes NN across the critical point.
+The critical field is located at $h_{i}=1$ for $\alpha=0.25$, and approximately at $h_{i}\approx 0.75$ for $\alpha=1.2$.
+To better understand optimization pathways, the behavior of thermodynamic efficiency $\eta$ and work per spin $W/N$ is analyzed as a function of the transverse field $h_{i}$ for different $N$ in Fig.
+7.a-b.
+For small $N$, high values of $\eta$ are observed at low $h_{i}$, despite a higher $W/N$ compared to larger systems.
+Near the critical point, $h_{i}=1$, short chains experience a sharp drop in efficiency, whereas longer chains exhibit a more gradual transition with a marked increase in $W/N$, displaying a more linear dependence on $N$.
+This contrasts with the refrigerator mode in Fig.
+7.c-d, where short chains yield a higher $\eta_{R}$, while longer chains extract more heat.
+Post-phase transition, longer chains predominate due to dominant short-range correlations, which reduce sensitivity to phase transitions and thermal fluctuations.
+In summary, this work has investigated the performance of quantum heat engines and refrigerators by analyzing the role of interaction range—encoded in the parameter $\alpha$—across different thermodynamic regimes.
 III Quantum heat engine and Refrigerator
@@
 Next, we will conduct a similar study on a quantum heat engine.
+Figure 10: Work per spin $W/N$ and efficiency $\eta$ as functions of the initial transverse field $h_{i}$ for different system size.
+For quenches performed in the paramagnetic phase, i.e., the critical field is located at $h_{i}=1$ for $\alpha=0.55$, and approximately at $h_{i}\approx 0.75$ for $\alpha=1.2$.
+for $h_{i}>h_{\text{crit}}$, a distinct structure becomes apparent.
+In particular, the first peak, which becomes more pronounced as $T_{c}$ increases, exhibits a superextensive scaling of the work with the system size $N$, suggesting a potential influence of quantum critical behavior, same analysis as in the top panels, now applied to the efficiency $\eta$ of the heat engine.
+The qualitative features remain consistent, including the dependence on NN near the critical peak.
 III.1 Quantum Refrigerator
-the refrigeration mode (R) is predominant when the thermal gap between the two reservoirs is small Fig.3.
-We will examine two values of $\alpha$: the first promotes long-range correlations ($\alpha=0.25$), and the second favors short-range correlations ($\alpha=1.2$) .
-We will also investigate the behavior of thermodynamic quantities such as efficacy, $\frac{Q_{c}}{N}$, and the scaling factor per spin, $\frac{\Pi}{N}$
-We notice in Fig.7($a$-$b$) that $Q_{c}/N$ for both values of $\alpha$ is of the same magnitude , for $\alpha=0.25$, small chains perform slightly better at low fields .
-We observe that it peaks just before the critical point $h_{i}=1$ and then slightly declines before and after the phase transition.
-It stabilizes after $h_{i}=2$ The contour lines of $\frac{Q_{c}}{N}$ for small $h_{i}$ are very tightly packed, particularly for short chains, showing greater variability and a more significant decrease compared to longer chains, which have less dense contour lines.
-It is observed that the contour lines indicate a slower and more stable decrease for all $N$ after $h_{i}=1.5$.
-In the case of $\alpha=1.2$, long chains outperform short ones before and around the critical field.
-In the critical field, $h_{\text{critique}}=0.7$, we observe a pronounced drop, additionally, we note that short-range correlations lead to a steeper and more abrupt decline compared to long-range correlations.We notice that for $Q_{c}/N$, the contour lines are less tight, with significant values for large $N$.
-Unlike the previous case, there is a more substantial decrease for long and intermediate chains, while the decrease is slower for short chains.
-A stabilization occurs for all chain sizes starting from $h_{i}=1.5$.
-Figure 7: Contour plot of $Q_{c}$ and $\eta_{R}$ as a function of $h_{i}$ and N.
-Figure 8: COP of the refrigerator versus $h_{i}$, for different system sizes $N$ across the critical point.
-Figure 9: Scaling factor per spin $\Pi_{R}/N$ as a function of $h_{i}$ for different N.
-The plot of COP as a function of $h_{i}$ Fig.8($a$) for different $N$ at $\alpha=0.25$ for $\frac{Q_{c}}{|Q_{c}+Q_{h}|}$ shows better efficiency for small chains (10 to 40), with a faster growth than larger chains, particularly near the critical field $h_{i}=1$.
-There is a distinct difference in behavior between large and small $N$, as in the post-critical region there is a sharp drop for small chains (10 to 40), with stabilization occurring from $h_{i}=1.8$.
-For intermediate and large chains, we observe a slight increase before the critical magnetic field, which becomes more pronounced after the phase transition, Starting from $h_{i}=1.5$, a shift in efficiency is observed: larger chains exhibit greater efficiency compared to intermediate and shorter chains.
-This highlights a stronger size dependence after the phase transition, particularly for higher values of $h_{i}$.
-Beyond the critical point, small chains struggle to recover their refrigeration efficiency, with $Q_{c}/N$ remaining low and showing no significant improvement as $h_{i}$ increases.
-This highlights the difficulty for smaller systems to adapt to the post-transition regime, resulting in a diminished capacity to extract heat effectively.
-In contrast, larger chains demonstrate a partial recovery in $Q_{c}/N$ as $h_{i}$ exceeds 1.5, suggesting their greater adaptability and resilience in the altered dynamics of the post-critical regime.
-While they do not reach the efficiency levels observed before the transition, large systems still maintain some level of refrigeration capability, indicating their robustness in maintaining performance despite the phase change.
-The contour lines Fig.7($c$) show that they are low for small $h_{i}$, particularly for long chains.
-As $h_{i}$ increases (with intermediate magnetic fields $h_{i}$ from 0.5 to 1.5), the lines become concave and expand as $N$ decreases, centered around the critical field where they reach a maximum for short chains before dropping sharply after higher $h_{i}$.
-It is noted that for strong magnetic fields, the efficiency of long chains increases gradually.
-The figure Fig.8($b$) showing the efficiency for $\alpha=1.2$ as a function of $h_{i}$ and $N$ demonstrates that short chains are more efficient than long chains, similarly to previous cases.
-For small $h_{i}$, there is a slight and similar increase for both long and short chains.
-As the critical field $h_{i}=0.7$ is approached, the slope increases, more pronounced for long chains, which rise rapidly to approach the efficiency of short chains for intermediate fields.
-For higher magnetic fields, there is a slower increase for long chains, with stabilization around $h_{i}=2$, and a faster increase for short chains.
-For large chains, saturation is observed for all values of $h_{i}$, and the increase in COP is more significant for short chains, indicating that they undergo a more gradual transition and can maintain effective refrigeration for all values of $h_{i}$.
-The efficiency contour lines are less tight compared to the previous case, showing an increase in efficiency as $h_{i}$ rises for all $N$, with a leveling of the efficiency near the critical field $h_{i}=0.7$.
-Efficiency increases in the intermediate field range, and the contour lines become vertical.
-For stronger fields, starting from $h_{i}=1.7$, the lines curve and become more concave, less uniform, and more efficient for short chains than for long ones, with a peak for short chains ($N=10$ to $20$) at $h_{i}=2.5$ (see Fig.7($d$)).
-The scaling factor per spin, $\Pi_{R}/N$, is influenced by the parameter $\alpha$, unlike $Q_{c}/N$, where the values are not of the same order.
-An increase in $\alpha$ amplifies the fluctuations in our quantum refrigerator.
-It is also important to note that increasing $\alpha$ brings the temperatures of the reservoirs closer together.
-For $\alpha=0.25$, there is a pronounced peak at low $h_{i}$ values, especially for very long chains, which diminishes as $h_{i}$ increases.
-This decline is more significant for shorter chains near the phase transition at $h_{i}=1$, where the slope decreases and stabilizes performance for larger $h_{i}$ values.
-At $h_{i}=2.25$, the performance of long and short chains converges, demonstrating a lack of size dependence for higher $h_{i}$ values.For $\alpha=1.2$, before the paramagnetic phase, performance stabilizes with a slight decline as $h_{i}$ increases across all chains.
-At the critical field $h_{i}=0.7$, there is a sharp drop in the scaling factor, signaling underperformance and instability, particularly for longer chains, which remain more efficient than smaller $N$ but exhibit a steeper decline.
-As $h_{i}$ increases, there is a modest decrease followed by stabilization (see Fig.(9)).
-Overall, the system’s behavior strongly depends on the chain size and the correlation regime, with each configuration offering specific advantages depending on the magnetic field.
+A thermodynamic advantage is expected in the vicinity of the phase transition as highlighted in [57], in finite-temperature systems near criticality, enhanced heat extraction is anticipated around the critical point.
+This enhancement arises from the divergence of the specific heat [58], which allows substantial heat exchange even under small temperature gradients.
+A comparable reasoning applies to our magnetic system: as the quantum critical point is approached, the magnetic susceptibility—defined as the derivative of magnetization with respect to the external field—diverges.
+Consequently, the magnetization becomes highly responsive to even minor variations in the field.
+Since the work output due to a magnetic field change is proportional to the magnetization [59, 60], this leads us to expect an increase in work extraction in the vicinity of the quantum critical point , similarly, improved heat extraction is expected in the refrigerator mode.
+Our findings, as depicted in Fig.
+8.a-b, demonstrate that for both values of $\alpha$, an increase in the interaction range enhances the strength of spin correlations and mitigates perturbations induced by neighboring particles, particularly in the ferromagnetic phase ($h_{i}<h_{c}$).
+This fosters coherent spin alignments, thereby improving efficiency.
+We observe superior heat extraction per spin, $Q_{c}/N$, for long-range interactions, which facilitate more effective energy transfer among nearest neighbors compared to short-range interactions that propagate more locally, thus minimizing $Q_{c}/N$.
+Long-range interactions lead to a smoother and more gradual decline in $Q_{c}/N$ beyond the critical transition, especially for longer chains, which exhibit greater stability and resilience.
+Conversely, short-range interactions result in a sharper drop around the critical point, attributable to the localized nature of energy diffusion compared to the more extended correlations observed for smaller $\alpha$.
+Drawing on the evidence provided by [14], the authors showed that taxonomy separates three regimes by $\alpha$.
+At $\alpha=0.2$ (very long range, effectively near all-to-all) the system belongs to the non-additive, strong-long-range sector: without Kac rescaling the interaction energy grows super-extensively; collective, mean-field–like correlations and long-lived quasi-stationary states appear, relaxation is anomalously slow, and pronounced finite-$N$ peaks or apparent saturations dominate observables.
+By contrast, at $\alpha=1.2$ the model lies in the crossover from weak long-range to short-range behaviour, where energy is extensive, critical exponents and finite-size corrections become largely $\alpha$-independent, and observables approach conventional short-range scaling smoothly.
+The analysis of the coefficient of performance (COP) corroborates the previously elucidated behavior in Fig.
+9.a-b, where an increase in $\alpha$ leads to an enhanced COP, particularly for short chains, which preferentially favor the refrigeration mode.
+In contrast, long-range interactions, under the influence of neighboring particles, amplify perturbations and spin disorder, a phenomenon that is accentuated in the paramagnetic phase for longer chains, thereby fueling their performance.
+The contour plots of $Q_{c}/N$ (Fig.
+8.a-b) and COP (Fig.
+8.c-d) highlight size-dependent dynamics.
+As $\alpha$ increases, contour lines become densely packed near the transition for small $N$, particularly for long chains, where the behavior is less smooth and more abrupt due to the nature of interactions, which are not as diffusive as long range interaction.
+Figure 11: Contour plot of the work per spin $W/N$ and $\eta$ as a function of $h_{i}$ for different sizes N , for different values of interaction range $\alpha$ .
+Overall, the observed deviations for small values of $N$ are indeed expected and can be attributed to well-documented finite-size effects, including discretized energy levels, boundary contributions, and the lack of a thermodynamic limit, cause fluctuations and non-extensive behaviors in quantities like heat and work per spin in small spin chains [61, 62].
+Quantum coherence and entanglement further influence the performance of quantum thermal machines at small $N$, leading to non-classical thermodynamic signatures.
+As $N$ increases, these effects diminish, and observables like work and heat per spin align with the extensivity principle of thermodynamics, consistent with known physics of finite-size systems [63, 64].
+The system’s behavior strongly depends on the chain size and the correlation regime, with each configuration offering specific advantages depending on the magnetic field.
 The performance per spin is better in the ferromagnetic phase, with efficiency varying depending on the type of correlation $\alpha$.
 III.2 Quantum Heat Engine
-We will see in this section the evolution of the work per spin ($W/N$) and efficiency ($\eta$) as a function of the initial transverse field ($h_{i}$) for different values of $\alpha$, $\beta_{h}$, and $\beta_{c}$, analyzing the effects of the phase transition and the system’s behavior for small and large chains ($N$).
-Each graph in Fig.(10) highlights how interactions and thermal conditions influence the system’s performance.
-Figure 10: Work per spin $W/N$ as a function of $h_{i}$ for different N.
-In Fig.10($a$), with $\alpha=0.9$, $\beta_{h}=0.5$, and $\beta_{c}=1.0$, it is observed that for large values of $N$ (e.g., $N>70$), the work per spin is higher in the region where $h_{i}\leq 1.0$.
-Beyond this point, the slope decreases, leading to a stabilization of the curve.
-The maximum work per spin, $W/N$, occurs around $h_{i}=0.25$, indicating a more effective capture of short-range correlations near the critical point.
-It is noteworthy that, across all regions, long chains exhibit a higher $W/N$ compared to short chains.
-After the critical point, $W/N$ declines rapidly, especially for large systems, signaling a significant shift associated with the phase transition.
-In contrast, for smaller systems, the maximum is less pronounced, and the work is lower, highlighting the dominance of thermal fluctuations and local correlations.
-When comparing the sizes of the chains, $W/N$ remains higher for long chains throughout the $h_{i}$ range from 0 to 2.5, demonstrating greater efficiency in capturing short-range correlations for longer chains.
-In Fig.10($b$), corresponding to $\alpha=1.2$, $\beta_{h}=0.2$, and $\beta_{c}=0.9$, the work per spin is similarly elevated for larger values of $N$, peaking around $h_{i}=0.35$, which aligns with the estimated critical point for $\alpha>1$ at $h=0.7$.
-Long chains are more adept at capturing short-range correlations than their shorter counterparts prior to the transition, thereby accounting for a higher $W/N$.
-After $h_{i}=0.7$, $W/N$ experiences a steep decline, occurring more rapidly than in the previous case, while the work per spin of long chains remains superior to that of short chains.
-This observation indicates that longer chains continue to effectively capture short-range correlations even subsequent to the phase transition.
-In contrast, shorter chains exhibit a more stable behavior, suggesting reduced sensitivity to fluctuations in $h_{i}$.
-The Fig.10($c$), with $\alpha=0.55$, $\beta_{h}=0.2$, and $\beta_{c}=0.667$, shorter chains ($N<30$) exhibit a higher $W/N$ for $h_{i}$ ranging from 0 up to the critical point at $h=1.0$, where a transition occurs, allowing long chains to become more efficient following the phase transition; this alteration does not apply to short-range correlations.
-The peak of $W/N$ is observed around $h_{i}=0.25$, indicating that smaller chains are more effective in capturing intermediate-range correlations prior to the phase transition.
-However, after $h_{i}=1$, longer chains ($N>70$) surpass shorter chains in terms of $W/N$, reflecting a superior ability to capture long-range correlations.
-This implies that for $\alpha=0.55$, the phase transition significantly alters the behavior of long chains post $h_{i}=1$, while shorter chains maintain their efficiency before this transition.
-Fig.10($d$), with $\alpha=0.2$, $\beta_{h}=0.6$, and $\beta_{c}=2.0$, exhibits behavior akin to the previous case, with the performance of long chains surpassing that of shorter chains before the critical point at $h=1$, particularly near $h=0.4$.
-For shorter chains ($N<30$), $W/N$ is elevated for $h_{i}\leq 0.4$, peaking around $h_{i}=0.8$ for longer chains.
-Beyond this threshold, $W/N$ declines rapidly, indicating a loss of efficiency in the smaller chains.
-In contrast, after the critical field, starting from $h_{i}=1$, there is a significant drop in $W/N$.
-This illustrates that long-range correlations in long chains at the phase transition accelerate the decline of $W/N$.
-The contour lines of the work per spin $W/N$ presented in Fig.11 show behavior that depends on $\alpha$ and the thermal gap.
-When $\alpha$ is small and favors long-range correlations, there are two peaks: one for short chains and the other for long chains unlike $\alpha=0.25$, we observe a peak before the transition (for small $N$) and another at the critical field (for large $N$).
-When $\alpha=0.55$, there is a peak before the critical field and another after the phase transition, which is farther apart with a lower $W/N$ at the critical field compared to the previous case.
-When short-range correlations are favored ($\alpha=0.9$ and $1.2$), we observe a single peak before the critical field for long chains.
-The only difference between the two cases lies in the variability and a smaller decrease in $W/N$ before the phase transition for $\alpha=1.2$, with more vertical contour lines.
-This detailed analysis presents the behavior of the efficiency $W/Q_{h}$ and the work per spin $W/N$ as a function of $h_{i}$ for different values of $\alpha$, $\beta_{h}$, and $\beta_{c}$.
-Figure 11: Contour plot of Work per spin W/N and efficient $\eta$ as a function of $h_{i}$ and N.
-With $\alpha=0.9$, $\beta_{h}=0.5$, and $\beta_{c}=1.0$, the efficiency $W/Q_{h}$ shows a gradual decline as $h_{i}$ increases.
-The phase transition, occurring at $h=1$, corresponds to the critical point $h_{c}$ for $\alpha\leq 1$.
-Prior to this transition, longer chains (with lower $N$) display greater efficiency compared to shorter chains.
-However, near the critical point, the efficiency of longer chains aligns closely with that of shorter ones.
-Beyond $h=1.25$, the disparity between short and long chains becomes more apparent, accompanied by a marked decrease in efficiency for both.
-In the second figure, with $\alpha=1.2$, $\beta_{h}=0.2$, and $\beta_{c}=0.9$, we observe a pattern reminiscent of the case with $\alpha=0.9$, where short-range correlations are predominant.
-The critical field is located at $h_{i}=0.7$.
-Prior to the phase transition, there is a slight reduction in efficiency for chains of all sizes, with large $N$ chains exhibiting greater efficiency.
-At the critical point, a distinctive behavior emerges with a sharp drop in efficiency, resulting in a significant decline across all chains, leading to an identical efficiency independent of chain length in the post-critical region.
-Efficiency remains relatively high for small values of $h_{i}$ (between 0 and 0.5), reaching a peak at very low fields.
-Large chains (with higher $N$) demonstrate superior efficiency compared to smaller chains across the entire range of $h_{i}$, both before and after the critical field.
-This stability in larger chains following the transition highlights a regime where short-range correlations prevail, maintaining efficiency despite variations in the transverse field $h_{i}$.
-In the third figure, with $\alpha=0.55$, $\beta_{h}=0.2$, and $\beta_{c}=0.667$, which favors intermediate range correlations, the efficiency $W/Q_{h}$ exhibits a gradual decline as $h_{i}$ increases, with nearly identical efficiency for large $N$ values, indicating pre-saturation for long chains in the range $h_{i}=0$ to 0.5, just before the phase transition.
-Around the critical field $h_{c}=1$ for $\alpha\leq 1$, a stabilization is noticeable, where efficiency remains almost unchanged.
-Beyond $h_{i}=1.5$, a slight reduction in efficiency is observed for all chain lengths.
-The curve’s behavior suggests heightened sensitivity of the system to changes in $h_{i}$ near the critical point.
-The Fig.12($a$), with $\alpha=0.2$, $\beta_{h}=0.6$, and $\beta_{c}=2.0$, the efficiency is initially high but steadily declines as $h_{i}$ increases, reaching the critical point at $h_{i}=1$.
-This behavior reflects a phase transition, where the system’s sensitivity to the transverse field $h_{i}$ is notably heightened for small values of $\alpha$.
-The slope increases and the decline becomes more rapid for larger $h_{i}$, emphasizing the influence of long-range correlations in stronger fields.
-Post-transition, shorter chains demonstrate greater stability with fewer fluctuations compared to longer chains.
-For large $N$, the efficiency becomes more uniform, exhibiting faster saturation, more pronounced than in the case of $\alpha=0.55$, though this saturation occurs in the post-transition region, accentuated by long-range correlation effects.
-In stronger fields, such as $h_{i}=2$, intermediate-length chains show superior efficiency, indicating a shift in the system’s operational regime where long-range correlations become less effective for longer chains (see Fig.12).
-Figure 12: Efficiency $\eta$ as a function of $h_{i}$ for different $N$.
-The contour lines of efficiency $\eta$ show a dependence on $\alpha$.
-Similar to $W/N$, there are two distinct behaviors.
-For small $\alpha$, we observe a peak for long chains at very low $h_{i}$, with the contour lines becoming more horizontal for medium and large $N$, then more vertical for long chains, which is less noticeable for $\alpha=0.55$.
-For larger $\alpha$, there is a clear peak for medium and large chains at low $h_{i}$, with vertical contour lines where efficiency increases horizontally.
-This effect is more pronounced for $\alpha=1.2$, which shows slower variability compared to $\alpha=0.9$ (Fig.11(e)-(h)).
-Unlike work, the efficiency for all the parameters previously mentioned is dependent on $N$.
-To increase efficiency for fixed parameters, it is sufficient to increase the number of particles, which is not always the case for $W/N$.
-We observe that the efficiency ($\eta$) and the work per spin ($W/N$) exhibit different behaviors, as the work per spin is not proportional to the chain size for small and intermediate values of $\alpha$.
-In contrast, the efficiency depends directly on the size of the chain: the larger $N$ becomes, the greater the efficiency ($\eta$).
-This discrepancy can be explained by the nature of the correlations and fluctuations induced by the temperature difference between the reservoirs.
-For larger values of $\alpha$ (e.g., 0.9 and 1.2), we observe that when the temperature difference ($\Delta T$) is significant (as with $\alpha=1.2$), the system achieves a higher work per spin ($W/N$) compared to $\alpha=0.9$, but with a lower efficiency.
-This can be attributed to thermal fluctuations and agitation, which are reduced when $\Delta T$ is smaller.
-The same interpretation applies to long-range and intermediate-range correlations, where the effect of temperature fluctuations diminishes with a smaller temperature gradient, leading to different performances in terms of both work and efficiency.
+We will see in this section the evolution of the work per spin ($W/N$) and efficiency ($\eta$) as a function of the initial transverse field ($h_{i}$) for different values of $\alpha$, $T_{h}$, and $T_{c}$, analyzing the effects of the phase transition and the system’s behavior for small and large chains ($N$).
+Each graph in Fig.11 highlights how interactions and thermal conditions influence the system’s performance.
+Fig 10.a-b elucidates the behavior of the work per spin, $W/N$, as a function of varying $\alpha$, $T_{h}$, and $T_{c}$, highlighting the influence of chain length $N$, transverse field $h_{i}$, and differing interaction ranges.
+Increasing $\alpha$ reduces the influence of neighboring particles while enhancing localized correlations, thereby promoting stronger spin alignment in the ferromagnetic phase ($h_{i}<h_{c}$).
+As illustrated in Fig 10-b, elevating $\alpha$ generally leads to improved work extraction per spin, owing to diminished perturbations from proximate neighbors, resulting in a more linear behavior for large $N$.
+This contrasts with the case of $\alpha=0.55$, where more pronounced perturbative effects from neighboring particles intensify with increasing $N$, leading to a rapid misalignment of spins and consequently higher work extraction for smaller $N$.
+In contrast, the paramagnetic phase ($h_{i}>h_{c}$), characterized by disordered spins, leverages these neighboring perturbations as a driving force, significantly enhancing work extraction, particularly for longer chains.
+Regarding the efficiency, $W/Q_{h}$, Figure 10c-d highlights distinct regimes where, in contrast to the work per spin, an increase in $\alpha$ diminishes the cycle’s efficiency.
+This is attributed to $Q_{h}$, the heat absorbed from the hot reservoir, which is more effectively catalyzed and absorbed under long-range interactions.
+Conversely, at $\alpha=1.2$, the more localized diffusion imposed by short-range interactions restricts this process, leading to reduced efficiency.Complementing this analysis, Fig.
+11.a-b illustrates the contour lines of $W/N$, revealing dependencies on $\alpha$ and the thermal gap.
+For small $\alpha$, two peaks emerge for short and long chains in phases exhibiting distinct behaviors within the same interaction range, characterized by concentric circles centered around small $N$ in the ferromagnetic phase and large $N$ in the post-critical phase.
+For larger $\alpha$, a single peak at $\alpha=1.2$ dominates in the post-critical phase for large chains, where the locality of short-range interactions introduces a linear dependence before the phase transition, followed by a gradual decline toward higher field strengths.
+Figure 12: The scaling factor per spin $\Pi/N$ ( $\Pi_{R}/N$) as a function of $h_{i}$, for various system sizes NN and different interaction ranges characterized by $\alpha$.
+Regarding efficiency, $W/Q_{h}$, Fig.
+11.c-d highlights distinct regimes.
+For $\alpha=0.55$, efficiency decreases gradually, with a rapid vertical drop for long chains before the phase transition, in contrast to $\alpha=1.2$, which exhibits a more tempered decline and varies horizontally due to the linearity of short-range interactions.
 Thus, the relationship between efficiency and work per spin reflects the complex interplay of thermal fluctuations and correlations in the system, especially across different interaction ranges.
-III.3 Scaling factor of the quatum heat engine
-The scaling factor $\Pi=\frac{W}{\eta_{\text{Carnot}}-\eta}$ measures how effectively an engine produces work despite having an efficiency lower than the Carnot limit.
-A high scaling factor means that the system is optimized to extract significant work, even in non-ideal conditions.
-In quantum engines, a higher $\Pi$ reflects more efficient energy conversion through quantum effects like correlations and superposition.
-It also suggests lower energy losses, allowing the system to approach Carnot efficiency.
-Overall, a high scaling factor indicates that the system is productive, efficient, and resilient to fluctuations.
-To observe the impact and influence of correlations without considering the size of the chain, we will normalize by dividing by $N$.We will discussed this in more detail in the appendix Appendix D.
-The corrected interpretation of the provided figures focuses on the scaling factor $\Pi$, defined as
+III.3 Scaling factor of the quantum heat engine and refrigerator
+The scaling factor $\Pi$ measures how effectively an engine produces work despite having an efficiency lower than the Carnot limit [65].
+Furthermore , an elevated $\Pi$ plays a pivotal role in criticality and phase transitions, where it indicates enhanced system stability against critical fluctuations and the capacity to sustain optimal performance during such transitions.
+Consequently, a high scaling factor reflects increased efficiency, resilience to perturbations, and optimal utilization of energy resources, even though the absolute efficiency remains below the theoretical maximum.To observe the impact and influence of correlations without considering the size of the chain, we will normalize by dividing by $N$ , defined as :
 $\Pi/N=\frac{W}{(\eta_{\text{Carnot}}-\frac{W}{Q_{h}})}/N,$ (22)
-where $W$ is the work done by the system, $\eta_{\text{Carnot}}=1-\frac{\beta_{h}}{\beta_{c}}$ is the Carnot efficiency, and $N$ is the number of particles in the system .
-The figures 13 illustrate $\Pi/N$ as a function of the transverse magnetic field $h_{i}$, highlighting the behavior of the system for varying values of $N$, ranging from 10 to 100, for different configurations of $\alpha$, $\beta_{h}$, and $\beta_{c}$.
-Fig.13($a$) $\alpha=0.9$, $\beta_{h}=0.5$ and $\beta_{c}=1.0$, $\Pi/N$ starts at high values for small $h_{i}$ and decreases significantly near the critical magnetic field $h_{i}=1$, indicating a strong phase transition.
-Large chains ($N>70$) show a sharp decrease in $\Pi/N$, while small chains show a more gradual decline.
-After $h_{i}=1$, $\Pi/N$ for large chains remains stable but lower compared to small $h_{i}$ values, indicating a more consistent capture of long-range effects.
-$\Pi/N$ drops significantly after the transition, reflecting their inability to maintain high performance in the long-range regime.
-In , corresponding to $\alpha=1.2$, Fig.13($b$)$\beta_{h}=0.2$ and $\beta_{c}=0.9$, $\Pi/N$ is high for large values of $N$ (for example, $N>70$) in the region where $h_{i}\leq 0.7$.
-The maximum occurs around $h_{i}=0.3$, reflecting better adaptation to critical fluctuations near the estimated transition point for $\alpha>1$.
-Beyond the critical point $h_{i}=0.7$, there is a sharp drop in the system’s performance, with a less steep slope for small chains.
-After $h_{i}=1$, $\Pi/N$ decreases for large chains but remains higher compared to small chains ($N<30$), indicating that large chains are more effective in maintaining high performance before the transition but experience a more intense drop in performance than small chains, where long-range correlations dominate.
-Large chains show a rapid decrease in $\Pi/N$, reflecting their limited capacity to harness critical effects.
-At $\alpha=0.55$ Fig.13($c$), $\beta_{h}=0.2$ and $\beta_{c}=0.667$, shows that $\Pi/N$ is initially high, especially for small chains.
-Where long-range correlations are moderate before the critical field $h=1$, in the ferromagnetic phase, they force the particles to align better than in long chains, thus ensuring better $\Pi/N$ performance.
-After the transition phase $h_{i}=1$, large chains dominate in terms of $\Pi/N$, reflecting their superior performance in exploiting intermediate correlations in the ferromagnetic phase at higher field strength.
-This behavior indicates that large chains are more adaptable and efficient after the transition, whereas small chains are more efficient in the ferromagnetic phase, before the phase transition.
-In Fig.13($d$), with $\alpha=0.2$, $\beta_{h}=0.6$ and $\beta_{c}=2.0$, where the interactions are the longest, we observe that before $h_{i}=0.25$, the short chains perform better, and there is a shift at this point towards the dominance of long chains.
-$\Pi/N$ peaks at $h_{i}=0.6$ before the critical point but decreases rapidly beyond this point, showing poor performance in the post-transition regime.
-We also observe that $\Pi/N$ decreases slowly after $h_{i}=1$, indicating an effective capture of long-range correlations.
-This highlights the dominance of long-range effects for large chains before the phase transition.
-It is also important to note that, as in the previous section, there is a slight shift between long and short chains post-phase transition, as the short chains stabilize more quickly because the fluctuations and thermal perturbations, amplified by long-range correlations in the paramagnetic phase.
-Figure 13: Contour plot of the scaling factor per spin $\Pi_{R}/N$ as a function of $h_{i}$ and N.
-The impact of $\alpha$, $\beta_{h}$, and $\beta_{c}$ on the system is significant.
-For high $\alpha$ values (e.g., $\alpha=1.2$), short-range correlations dominate, leading to a smoother transition and more stable $\Pi/N$.
-Large chains benefit significantly from short-range interactions before $h_{c}$.
-For low $\alpha$ before critical field long-range effects are pronounced, leading to sharper transitions in $\Pi/N$ and small chains are more effective in these regimes, showing larger values of $\Pi/N$.
-A higher thermal gradient enhances the sensitivity of $\Pi/N$ to variations in $h_{i}$.
-For example, in the third figure, the large difference between $\beta_{c}=2.0$ and $\beta_{h}=0.6$ causes a significant shift in $\Pi/N$ post-transition.faor a small thermal gap (e.g., $\beta_{c}=1.0,\beta_{h}=0.5$ in the second figure) result in a smoother decline of $\Pi/N$, with less abrupt changes across the critical point but with a lower performance than the others.
-Before the critical point ($h_{i}<1$), small chains exhibit higher $\Pi/N$, reflecting their ability to efficiently capture long-range correlation in weak field $h_{i}$ , while large chains have a lower $\Pi/N$ but show a better performance than the shorter chain for higher $h_{i}$ special in the vicinity of critical point and after phase transition, indicating the growing influence of long-range correlations after the critical point ($h_{i}>1$), large chains outperform small chains, with a higher $\Pi/N$ due to the dominance of long-range correlations.
-Small chains show a significant decline in $\Pi/N$, reflecting reduced efficiency in capturing long-range effects.
-At the critical point ($h_{i}=1$), $\Pi$ peaks for large chains as the system approaches Carnot efficiency, reflecting enhanced work output and system efficiency.
-For small chains, the peak is less pronounced, indicating their limited capacity to utilize critical phenomena effectively.
-Refer to captionRefer to caption
-; Refer to captionRefer to caption
-Figure 14: The evolution of the maxima as a function of $N$ (the number of particles) for the diamagnetic peak (in red) and the paramagnetic peak (in blue) of $\Pi/N$ and $\Pi_{R}/N$, with a comparison to the critical field (in red).
-For the case of short-range correlations, we observe an underperformance after the phase transition, but with a stronger dependence on system size, as longer chains prove more efficient both before and after the critical point.
-It is also noted that for large $\alpha$ values, short-range correlations dominate, and the system’s efficiency is enhanced by the thermal gap, which promotes short thermal correlations and better protects them from thermal fluctuations.
+where $W$ is the work done by the system, $\eta_{\text{Carnot}}=1-\frac{T_{c}}{T_{h}}$ is the Carnot efficiency, and $N$ is the number of particles in the system in Fig.
+12 .
+These findings underscore the interplay of chain size, field strength, and interaction range in dictating system performance across phase transitions.
+Physically, $\Pi/N$ reflects the system’s capacity to harness quantum correlations for work output, modulated by the interplay between $\alpha$ and the thermal gradient.
+Near criticality in finite-temperature systems, enhanced heat extraction is anticipated due to the divergence of specific heat, allowing significant heat exchange with minimal temperature gradients.
+Similarly, in this magnetic system, the magnetic susceptibility—diverges near the quantum critical point, making magnetization highly responsive to small field changes.
+Since work output from a magnetic field variation is proportional to magnetization , this heightened sensitivity boosts work extraction at criticality.An increase in $\alpha$ in Fig.
+12-b enhances the strength of spin-spin correlations and reduces perturbations induced by neighboring particles, particularly in the ferromagnetic phase ($h_{i}<h_{c}$), promoting coherent spin alignments and elevating $\Pi/N$ by maximizing work extraction.
+Conversely, lower $\alpha$ values in Fig.
+12-a lead to long-range interactions with faster decay, increasing scattering and broadening the phase transition into a wider critical region .
+This renders short chains more stable before the phase transition.
+This diminishes the magnitude of $\Pi/N$ in the ferromagnetic phase due to weaker correlations but mitigates the impact of critical fluctuations, resulting in a smoother decline of $\Pi/N$ across $h_{c}$, particularly in smaller systems less sensitive to long-range effects .
+Let’s move on to the refrigerator mode, we operate with $\Pi_{R}/N$,
+$\Pi_{R}/N=\frac{Q_{c}}{(\eta_{\text{RCarnot}}-\frac{Q_{c}}{W})}/N,$ (23)
+where $Q_{c}$ denote the heat exchanged during the thermalisation with the cold bath done by the system, $\eta_{\text{RCarnot}}=\frac{1}{T_{h}/T_{c}-1}$ is the Carnot refrigerator efficiency , which reflects the system’s capacity to efficiently extract heat by leveraging spin correlations, modulated by $\alpha$ and the thermal gradient.
+Similarly to the engine mode, an increase in $\alpha$ influences the strength of correlations with neighboring spins (Fig.
+12-c) .
+Figure 13: The evolution of the maxima as a function of NN (the number of particles) for the ferromagnetic peak (in red) and the paramagnetic peak (in blue) of $\Pi/N$ and $\Pi_{R}/N$, with a comparison to the critical field (in red).We will also analyze the peaks in the paramagnetic phase before the phase transition and the ferromagnetic phase after the transition, as well as their proportionality and dependence on N for the quantum heat engine and refrigerator mode by using a regression to calculate the exponent $a$ responsible for the growth rate of the scaling factor per spin, proportional to $N^{a}$.
+However, unlike the engine mode, heat extraction from the reservoir is more effective when $\alpha$ is lower.
+This is explained by the divergence of the specific heat near the phase transition, which diminishes as $\alpha$ increases, resulting in a reduced potential for heat extraction and thus providing greater stability for long-range interactions.
+In the presence of short-range correlations, post-transition performance declines, with a notable dependence on system size; longer chains demonstrate greater efficiency across all regimes.
+Additionally, for large $\alpha$, short-range correlations prevail, bolstered by the thermal gap, which mitigates thermal fluctuations and enhances system efficiency.
 The behavior of $\Pi/N$ as a function of $h_{i}$ demonstrates the intricate relationship between work, system efficiency, and particle number $N$.
 Large chains tend to dominate in regimes where long-range correlations are significant, especially beyond the critical point, whereas small chains are more effective in the short-range, pre-transition regime.
-The parameters $\alpha$, $\beta_{h}$, and $\beta_{c}$ play crucial roles in shaping this behavior, with their values determining the system’s sensitivity to field variations and the overall efficiency relative to Carnot.
-Understanding these dynamics is essential for optimizing quantum heat engines across different operating regimes.
 IV Investigating the role of thermal correlations and criticality
-In this section, we will examine the scaling factor $\Pi/N$, studying its impact on the system’s behavior for varying values of $N$ (ranging from 10 to 100) and different configurations of $\alpha$, $\beta_{h}$, and $\beta_{c}$.
+In this section, we will examine the scaling factor $\Pi/N$, studying its impact on the system’s behavior for varying values of $N$ (ranging from 10 to 100) and different configurations of $\alpha$, $T_{h}$, and $T_{c}$ .
 focusing on the behavior near the critical field, including the peaks observed in the paramagnetic and ferromagnetic phases.
@@
 We will then examine the behavior of the performance and its trend for small, intermediate, and long chains and the role of thermal gaps and correlations $\alpha$ , to observe the impact of the thermal reservoirs on our system.
-In the case where $\alpha=0.9$, with $\beta_{h}=0.5$ and $\beta_{c}=1.0$ Fig.14($a$), the obtained exponents are $a=0.4578$ for the critical field, and 0.5154 and 0.46 respectively for the ferromagnetic and paramagnetic phases , with a poor temperature difference between the reservoirs.
-The scaling factor per particle describes the performance per particle while accounting for the influence of others.
-We observe that the post-critical and critical exponents are almost identical, with a significant slope for small $N$ (10 to 40) and medium $N$ (40 to 70).
-This suggests that critical effects slow down the growth and stabilize it in the paramagnetic phase, especially for large $N$ (70 to 100).
-In the ferromagnetic phase, we see greater growth, with a slight decrease in slope for larger $N$, and the curve is higher than in the critical and paramagnetic cases.
-For $\alpha=1.2$, with $\beta_{h}=0.2$ and $\beta_{c}=0.9$ Fig.14($b$), the obtained exponents are $a=0.2355$ for the critical field, 0.2308 and 0.2355 for the ferromagnetic and paramagnetic phases , respectively.
-In this case, $\alpha>1$, which favors more short-range correlations, and the thermal gap is higher compared to the previous case.
-The exponents are smaller than for $\alpha=0.9$ and are very close to each other, with one of the curves (critical and post-critical) perfectly overlapping The paramagnetic phase exhibits slightly lower growth than the ferromagnetic phase, but both show faster growth with $N$.
-A similar behavior is observed for small to intermediate $N$, but a notable difference appears for large $N$, where the critical and paramagnetic curves show the onset of pre-saturation the slope increases only marginally between $N=90$ and $N=100$.
-This indicates that for larger $\alpha$, the system’s performance depends on its size, which significantly amplifies thermal fluctuations and correlations, potentially leading to a more optimal yield.
-For Fig.14($c$) $\alpha=0.55$, which favors intermediate correlations, with a high thermal gap $\beta_{h}=0.2$ and $\beta_{c}=0.667$ Before the critical field, we observe a curve that is higher than the critical field and the post-phase transition, but it shows a downward behavior indicating underperformance of intermediate chains, which slows down around $N=90$.
-Near $N=55$ In the ferromagnetic phase, the spins are better aligned, but increasing the number of particles makes the system more sensitive to fluctuations caused by the different types of correlations, which increase the noise and cause the decrease of the scaling factor per spin ($\alpha=-0.1035$).
-There is an overestimation of the scaling factor per spin of the post-critical field peaks ($\alpha=0.2432$), with an increase of the performance of long chains.
-At the critical point, there is a slight and slow increase for long chains ($\alpha=0.0504$), indicating a greater difficulty aligning the particles for intermediate correlations before and during the phase transition.
-In the case of $\alpha=0.2$, with $\beta_{h}=0.6$ and $\beta_{c}=2.0$, Fig.14($d$) the thermal gap is short, low values of $\alpha$ favor long-range correlations, indicating that the fields before the critical point are more efficient.
-For all chains, performance in the phase transition and after the critical field shows a slight decrease in intermediate N, while the curves for the critical and post-critical peaks demonstrate an increase and outperform of the high chains, which is proportional to $N$.A slight under performance is observed in long chains (from $70$ to $100$) before the phase transition.
-The scaling factor per spin in the critical fields and after the transition the curves are superimposed for intermediate and shorter chains.
-This indicates that in long correlation before the critical field there is a superior performance and better spin alignment, resulting in less dissipation and a higher efficiency, there is nonetheless slower growth ($\alpha=0.32826$) compared to postcritical fields ($\alpha=0.9434$) and at the phase transition ($\alpha=0.9413$).
-We will then move on to the refrigeration mode, which is more prevalent in regions with a small thermal gap.
-We will examine the influence of the parameter $\alpha$, considering both small and large values, which, respectively, promote long-range and short-range thermal correlations.
-In the case of $\alpha=0.25$, with $\beta_{h}=1.4$ and $\beta_{c}=1.5$, Fig.14($e$)there is a perfect superposition for intermediate and long chains, with a slight difference for small $N$ for the critical ($\alpha=1.6150$) and postcritical fields ($\alpha=1.2702$).
-We observe that the performance per particle of the post phase transition for all N is lower than the pre phase transition, which exhibits a significant increase for small chains and saturates at intermediate $N$ ($\alpha=0.5338$), then slightly decreases.
-In contrast, the critical field reduces the gap for larger chains.
-It is also noted that the performance before the phase transition is less dependent on the size, which is an effect of long-range correlations.
-For of $\alpha=1.2$, with $\beta_{h}=0.58$ and $\beta_{c}=0.6$, Fig.14($f$) there is a favoring of short-range correlations.
-Similarly, there is a perfect superposition between the critical fields and the post-critical peaks ($\alpha=0.4717$) and the the pics before the phase transition increase more slowly ($\alpha=0.4071$), however unlike the previous case they exhibit similar behavior with an increase as the chain size increases, but with a slightly smaller slope for the pre-critical case, which remains higher than the others regardless of $N$.
-This demonstrates that short-range correlations impose a chain-length dependency.
-The scaling factor $N^{\alpha}$ is crucial for understanding how the yield of a quantum thermal engine behaves as the system size increases.
 The goal is to identify conditions under which increasing $N$ (number of spins or system size) leads to a significant improvement in yield compared to the energy cost.
-When $\alpha>0$, it means that increasing $N$ results in a superlinear growth of cycle work and engine efficiency, which is a sought after goal to achieve better performance on a large scale.
-Near a critical point, energy fluctuations and heat capacity can diverge, leading to faster yield growth (higher value of $\alpha$).
-This is where critical engines can potentially achieve efficiency close to the Carnot point without sacrificing power.
-For lower values of $\alpha$ (typically far from the critical point), the yield per spin reaches saturation, as spin interactions are not coherent enough to benefit from increasing system size.
-In conclusion, the results show that the dynamics of thermal efficiency and work per spin depend strongly on the value of $\alpha$, which defines the range of interactions between spins.
-Temperatures ($\beta_{h}$ and $\beta_{c}$) also influence the nature of transitions, with marked differences in spin thermal sensitivity at different values of $h_{i}$.
-The observed critical points are generally in agreement with theoretical predictions, although some deviations suggest significant correlation effects in these regimes.
-We generally observe better performance when the thermal gaps are very large for $\alpha=1.2$ and $\alpha=0.55$.
-When short-range correlations are favored, we see a more linear critical behavior that depends more on $N$, particularly with the phase transition.
-For lower $\alpha$, there is less dependence on the chain length, which can experience a drop of $\pi/N$ for long-range drops ($\alpha=0.55$), or similar behavior before and after the phase transition for large $N$ ($\alpha=0.2$).
+When $\alpha>0$, it means that increasing $N$ results in a superlinear growth of cycle work and engine efficiency, to achieve better performance on a large scale.
+The parameter $\alpha$ governs the decay of spin interactions, transitioning from a long-range regime ($\alpha\leq 1$) to a short-range regime ($\alpha>1$).
+At a low $\alpha=0.2$ ($T_{h}=1.67$, $T_{c}=0.5$) , long-range interactions promote strong spin coherence in the ferromagnetic phase ($h_{i}<h_{c}$) for short chains, followed by a sharp decline for larger $N$, resulting in a negative ferromagnetic exponent (a=0.1035a=-0.1035) in Fig.
+13-a .
+This is attributed to a collapse of spin alignment due to heightened susceptibility to thermal and quantum perturbations.
+Longer chains exhibit modest growth, as long-range interactions enhance fluctuations across a greater number of spins, manifesting more prominently in larger systems with increased resilience post-phase transition.
+For a high $\alpha=1.2$ ($T_{h}=5.0$, $T_{c}=1.11$) Fig.
+13-b , short-range interactions yield small, consistent exponents ($a\approx 0.2355$) across all phases and all $N$, producing stable growth due to the rapid decay of perturbative long-range interactions.
+This restricts spin coherence, mitigating abrupt losses at $h_{c}$ and favoring the ferromagnetic phase.A pronounced amplification of the critical peak becomes apparent when examining the dependence of performance on system size.
+As previously discussed, the paramagnetic peak in $\Pi/N$ exhibits a standard linear growth behavior, i.e., $\Pi\sim N$, particularly in the case of long-range interactions characterized by the parameter $\alpha$.
+In contrast, the peak associated with the critical transition displays a markedly superlinear increase for short-range interactions, while it is suppressed in the presence of long-range, nonlocal interactions.Similarly, the refrigerator exhibits analogous behavior in Fig.
+13.c-d , resulting in an increased linearity of $\Pi_{R}/N$ in the ferromagnetic phase for $\alpha=1.2$, with a slope coefficient rising from 0.47 to 0.94(in Fig.
+13-d) .
+This indicates a size-dependent treatment across both phases, with a less pronounced drop compared to the engine mode.In systems with long-range interactions, paramagnetic peaks are enhanced and gradually decrease as the system size increases, eventually saturating to a finite value.
+This behavior contrasts with that observed for short-range interactions, where the maximum values of the critical peaks follow a clearly superlinear power-law scaling, well described by:
+$\Pi_{\text{crit}}^{(R)}/N\sim N^{\alpha},\quad(\alpha>0).$
 In conclusion, the performance of quantum thermal machines, whether as engines or refrigerators, is intricately linked to the scaling behavior governed by $\alpha$.
@@
 V Conclusion
-In this study, we investigated the performance of a long-range Ising chain quantum Otto engine and refrigerator under ideal conditions, focusing on the effects of system size $N$, the interaction range parameter $\alpha$, and the temperature ratios $\beta_{h}/\beta_{c}$.
+In this study, we investigated the performance of a long-range Ising chain quantum Otto engine and refrigerator under ideal conditions, focusing on the effects of system size $N$, the interaction range parameter $\alpha$, and the temperature ratios $T_{c}/T_{h}$.
 Our analysis demonstrates that thermal correlations significantly influence the operational efficiency of these quantum machines, particularly near the critical points of phase transitions.
@@
 For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the precritical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region.
-Similarly, for the refrigerator, long-range correlations are also favored in the post-critical field, especially in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$.
-We then fixed $\alpha$ and varied $N$ to investigate the influence of chain size on performance.
+Similarly, for the refrigerator, long-range correlations are also favored in the post-critical field, especially in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$.We then fixed $\alpha$ and varied $N$ to investigate the influence of chain size on performance.
 For the quantum refrigerator, our results indicate that the cooling efficiency is maximized for smaller system sizes before the phase transition.
@@
 This stability is attributed to the absence of non-adiabatic effects and the slow, quasi-static nature of the transformation.
-Additionally, the yield $\eta$ differs from that of the refrigerator, as it depends on the chain size for all values of $\alpha$.
-Overall, this study highlights the importance of distinguishing between the critical phase and the diamagnetic phase in understanding the performance of quantum heat engines and refrigerators.
-The ferromagnetic phase, characterized by significant fluctuations and superior performance, contrasts with the diamagnetic phase, which occurs beyond the critical point and exhibits greater stability but inferior performance.
+Additionally, the yield $\eta$ differs from that of the refrigerator, as it depends on the chain size for all values of $\alpha$.Overall, this study highlights the importance of distinguishing between the critical phase and the paramagnetic phase in understanding the performance of quantum heat engines and refrigerators.
+The ferromagnetic phase, characterized by significant fluctuations and superior performance, contrasts with the paraagnetic phase, which occurs beyond the critical point and exhibits greater stability but inferior performance.
 These findings underscore the need for a deep understanding of both static and dynamic conditions to effectively optimize the operation of quantum thermal machines.
-Appendix A Dynamic of the system
-A stationary state is reached in the long-time limit $t\to\infty$.
-In the following, we will characterize this stationary state through observables.
-We start with the adjoint version of our master equation, which can be written as:
-$\frac{dO_{H}(t)}{dt}=i\{H,O_{H}(t)\}_{-}+\sum_{n,k}\gamma_{n,k}\Big{(}[1-\zeta f_{n}(\omega_{k})]\\ \times[2b_{k}^{\dagger}O_{H}(t)b_{k}-\{b_{k}^{\dagger}b_{k},O_{H}(t)\}_{+}]\\ +f_{n}(\omega_{k})[2b_{k}O_{H}(t)b_{k}^{\dagger}-\{b_{k}b_{k}^{\dagger},O_{H}(t)\}_{+}]\Big{)},$ (9)
-where $O_{H}(t)$ denotes the Heisenberg form of a Schrödinger observable $O$.
-$\frac{d}{dt}\langle b_{k}^{\dagger}b_{k}\rangle$ $=-2\sum_{n}\gamma_{n,k}\langle b_{k}^{\dagger}b_{k}\rangle+2\sum_{n}\gamma_{n,k}f_{n}(\omega_{k}),$ (64a)
-$\frac{d}{dt}\langle b_{k}^{\dagger}b_{q}\rangle$ $=\Big{(}i(\tilde{\omega}_{k}-\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{k}^{\dagger}b_{q}\rangle,$ (64b)
-$\frac{d}{dt}\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle$ $=\Big{(}i(\tilde{\omega}_{k}+\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle,$ (64c)
-$\frac{d}{dt}\langle b_{q}b_{k}\rangle$ $=\Big{(}-i(\tilde{\omega}_{k}+\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{q}b_{k}\rangle.$ (64d)
-Developing the expressions, we find:
-$\langle b_{k}^{\dagger}b_{k}\rangle(t)$ $=\frac{\sum_{n}\gamma_{n,k}f_{n}(\omega_{k})}{\sum_{n}\gamma_{n,k}}\Big{(}1-\exp\big{(}-2\sum_{n}\gamma_{n,k}t\big{)}\Big{)}$
-$\quad+\langle b_{k}^{\dagger}b_{k}\rangle_{0}\exp\big{(}-2\sum_{n}\gamma_{n,k}t\big{)},$ (65a)
-$\langle b_{k}^{\dagger}b_{q}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}\rangle_{0}\exp\Big{(}i(\tilde{\omega}_{k}-\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65b)
-$\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{0}\exp\Big{(}i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65c)
-$\langle b_{q}b_{k}\rangle(t)$ $=\langle b_{q}b_{k}\rangle_{0}\exp\Big{(}-i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65d)
-where the subscript $\langle\cdot\rangle_{0}$ indicates the expectation value at time $t=0$. If $\sum_{n}\gamma_{n,k}\neq 0$, the limit $t\to\infty$ gives:
-$\langle b_{k}^{\dagger}b_{q}\rangle_{s}$ $=\delta_{kq}\frac{\sum_{n}\gamma_{n,k}f_{n}(\omega_{k})}{\sum_{n}\gamma_{n,k}},\quad\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{s}=\langle b_{q}b_{k}\rangle_{s}=0.$ (66)
-The equations (65) cease to be valid if, for some pair $(k,q)$, $\gamma_{n,k}=\gamma_{n,q}=0$ for all $n$. In that case, we find:
-$\langle b_{k}^{\dagger}b_{q}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}\rangle_{0}e^{i(\tilde{\omega}_{k}-\tilde{\omega}_{q})t},$ (67a)
-$\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{0}e^{i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t}.$ (67b)
-Appendix B Adiabatic Transformation
+Appendix A Long range chain Ising
+As a prototypical example of a many-body working substance with, possibly, long-range interactions, we consider a generic model of spinless fermions hopping across the $N$ sites of a linear chain in the presence of pairing interaction, and with a chemical potential $h$.
+Assuming periodic boundary conditions, the system Hamiltonian reads
+$H=$ $-\sum_{j=1}^{N}\sum_{r=1}^{N/2-1}\left[t_{r}c^{\dagger}_{j+r}c_{j}+\Delta_{r}c^{\dagger}_{j+r}c^{\dagger}_{j}+\text{h.c.}\right]$
+$-h\sum_{j=1}^{N}\left[1-2c^{\dagger}_{j}c_{j}\right],$ (24)
+where $c^{\dagger}_{j}$ and $c_{j}$ are creation and annihilation operators for fermions at site $j$, while $t_{r}$ and $\Delta_{r}$ are the hopping and pairing amplitudes, respectively.
+We choose their dependence on the intersite distance $r$ according to the power laws
+$t_{r}=\frac{1}{N_{\alpha}}\frac{J}{r^{\alpha}},\quad\Delta_{r}=\frac{1}{N_{\alpha}}\frac{J}{r^{\alpha}},$ (25)
+with the hopping exponent $\alpha_{1}>0$, the pairing exponent $\alpha_{2}>0$, and $N_{\alpha}=\sum_{r=1}^{N/2}r^{-\alpha}$ the Kac scaling factor, which guarantees extensivity of the energy in the case $\alpha_{i}<1$, with $i=1,2$.
+Hereafter, we set $J=\Delta=1$ as the energy scale and work in units of $\hbar=k_{B}=1$.
+In particular, in the short-range case with $\alpha\to\infty$, the relation becomes exact through the Jordan–Wigner mapping anad we descibes the ising transversed field model [66].
+Due to the translational invariant nature of the couplings, it is useful to write the Hamiltonian in terms of the momentum-space operators
+$\tilde{c}_{k}=e^{-i\frac{\pi}{4}}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{ikj}c_{j},$ (26)
+where $k=2\pi n/N$, with $n=-N/2+1,\dots,N/2$ (in the following we will drop the $\tilde{}$ on the $c_{k}$ unless it is ambiguous). Then we obtain
+$H=\sum_{k}\left[(h-t_{k})\left(c^{\dagger}_{k}c_{k}-c_{-k}c^{\dagger}_{-k}\right)+\Delta_{k}\left(c^{\dagger}_{k}c^{\dagger}_{-k}+c_{-k}c_{k}\right)\right],$ (27)
+where $t_{k}$ and $\Delta_{k}$ are the Fourier transforms of the hopping and pairing amplitudes, respectively, which in the thermodynamic limit may be written as
+$t_{k}$ $=\frac{\mathrm{Re}\left[\text{Li}_{\alpha_{1}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{1})},$ (28)
+$\Delta_{k}$ $=\frac{\mathrm{Im}\left[\text{Li}_{\alpha_{2}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{2})},$ (29)
+where $\text{Li}_{\alpha}(z)$ denotes the polylogarithm and $\zeta(\alpha)$ is the Riemann zeta function.
+We notice that the Hamiltonian in the Fourier space can be decomposed into the sum of single mode Hamiltonians, introducing $\Psi_{k}=(c_{k},c^{\dagger}_{-k})^{T}$
+$H=\sum_{k}\Psi^{\dagger}_{k}H_{k}\Psi_{k},$ (30)
+$H_{k}=(h-t_{k})\sigma^{z}_{k}+\Delta_{k}\sigma^{x}_{k},$ (31)
+[…]
+$u_{k}=\cos\frac{\theta_{k}}{2},\quad v_{k}=\sin\frac{\theta_{k}}{2},$ (32)
+where $\theta_{k}=\arctan\left[\Delta_{k}/(h-t_{k})\right]$, to obtain
+$H=\sum_{k}\omega_{k}(h)\left(\gamma^{\dagger}_{k}\gamma_{k}-\frac{1}{2}\right),$ (33)
+with the spectrum
+$\omega_{k}(h)=2\sqrt{(h-t_{k})^{2}+\Delta_{k}^{2}}.$ (34)
+Appendix B The Nambu formalism
+As we have seen, in the ordered case the Hamiltonian can be diagonalized by a Fourier transformation, reducing the problem to a collection of $2\times 2$ “pseudo-spin-1/2” problems, followed by a Bogoliubov transformation, .
+In the disordered case, we can proceed similarly, but we cannot reduce ourselves to $2\times 2$ problems in a simple way.
+By using the Nambu formalism, we define a column vector $\bm{\Psi}$ and its Hermitian conjugate row vector $\bm{\Psi}^{\dagger}$, each of length $2L$, by
+$\bm{\Psi}=\begin{pmatrix}\hat{c}_{1}\\ \vdots\\ \hat{c}_{L}\\ \hat{c}^{\dagger}_{1}\\ \vdots\\ \hat{c}^{\dagger}_{L}\end{pmatrix}=\begin{pmatrix}\hat{\mathbf{c}}\\ \hat{\mathbf{c}}^{\dagger}\end{pmatrix}$, (35)
+[…]
+Warning: Notice that the $\bm{\Psi}$ satisfy quite standard fermionic anti-commutation relations
+$\{\Psi_{j},\Psi^{\dagger}_{j^{\prime}}\}=\delta_{j,j^{\prime}},$ (36)
+for $j,j^{\prime}=1,\dots,2L$, except that $\{\Psi_{j},\Psi_{j+L}\}=1$ for all $j\leq L$, which brings about certain factors 2 in the Heisenberg’s equations of motion (see later).
+It is useful, for later purposes, to introduce the $2L\times 2L$ swap matrix $S$:
+$S=\begin{pmatrix}0_{L\times L}&1_{L\times L}\\ 1_{L\times L}&0_{L\times L}\end{pmatrix}$, (37)
+in terms of which $\bm{\Psi}^{\dagger}=(S\bm{\Psi})^{T}$.
+Consider now a general fermionic quadratic form
+$\hat{H}=\sum_{j,j^{\prime}}\left(A_{j^{\prime}j}\hat{c}^{\dagger}_{j^{\prime}}\hat{c}_{j}+A^{*}_{j^{\prime}j}\hat{c}^{\dagger}_{j}\hat{c}_{j^{\prime}}\right)+\sum_{j,j^{\prime}}\left(B_{j^{\prime}j}\hat{c}^{\dagger}_{j^{\prime}}\hat{c}^{\dagger}_{j}+B^{*}_{j^{\prime}j}\hat{c}_{j}\hat{c}_{j^{\prime}}\right),$ (38)
+where $A_{j^{\prime}j}=A^{*}_{jj^{\prime}}$, i.e., $A=A^{\dagger}$ is Hermitian, and $B_{jj^{\prime}}=-B_{j^{\prime}j}$, i.e., $B=-B^{T}$ is antisymmetric because $\hat{c}_{j}\hat{c}_{j^{\prime}}$ is antisymmetric under exchange of the two operators, and any symmetric part of $B$ would not contribute.
+It is simple to verify that $\hat{H}$ can be expressed in terms of $\bm{\Psi}$, omitting an irrelevant constant term $\operatorname{Tr}A$, as:
+$\hat{H}=\bm{\Psi}^{\dagger}H\bm{\Psi}=\begin{pmatrix}\hat{\mathbf{c}}^{\dagger},\hat{\mathbf{c}}\end{pmatrix}\begin{pmatrix}A&B\\ -B^{*}&-A^{*}\end{pmatrix}\begin{pmatrix}\hat{\mathbf{c}}\\ \hat{\mathbf{c}}^{\dagger}\end{pmatrix}$.
+(39)
+There is an intrinsic particle-hole symmetry in a fermionic Hamiltonian having this form.
+This symmetry, further discussed in Sec.
+6.1, is connected with the fact that the Hermitian $2L\times 2L$ matrix $H$ satisfies:
+$HS=-SH^{*}.$ (40)
+In the XY-Ising case, all couplings are real and we have two different fermionic Hamiltonians, one for each parity sector $p=0,1$, which we report here for convenience, using…
+Let us consider the eigenvalue problem for a general Hermitian $2L\times 2L$ matrix $H$
+$H\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}=\begin{pmatrix}A&B\\ -B^{*}&-A^{*}\end{pmatrix}\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}=\epsilon_{\mu}\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}$, (41)
+where $u,v$ are $L$-dimensional column vectors, composing the $2L$-dimensional column vector $\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}$, and the $\mu$ index refers to $\mu$-th eigenvector.
+By explicitly writing the previous system, we find the so-called Bogoliubov-de Gennes equations:
+$\begin{cases}Au_{\mu}+Bv_{\mu}=\epsilon_{\mu}u_{\mu},\\ -B^{*}u_{\mu}-A^{*}v_{\mu}=\epsilon_{\mu}v_{\mu}.\end{cases}$ (42)
+It is easy to verify that if $(u_{\mu},v_{\mu})^{T}$ is eigenvector with eigenvalue $\epsilon_{\mu}$, then $(v^{*}_{\mu},u^{*}_{\mu})^{T}$ is an eigenvector with eigenvalue $\epsilon_{\mu}$.
+We can organize the eigenvectors in a unitary (orthogonal, if the solutions are real) $2L\times 2L$ matrix
+$U=\begin{pmatrix}u_{1}&\cdots&u_{L}&v^{*}_{1}&\cdots&v^{*}_{L}\\ v_{1}&\cdots&v_{L}&u^{*}_{1}&\cdots&u^{*}_{L}\end{pmatrix}=\begin{pmatrix}U&V^{*}\\ V&U^{*}\end{pmatrix}$, (43)
+$U$ and $V$ being $L\times L$ matrices (real, as we can choose to be, in the Ising case) with the $u_{j}$ and $v_{j}$ as columns. As a consequence:
+$U^{\dagger}HU=$ $\begin{pmatrix}\epsilon_{1}&0&\cdots&0&0&0&\cdots&0\\ 0&\epsilon_{2}&\cdots&0&0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\epsilon_{L}&0&0&\cdots&0\\ 0&0&\cdots&0&-\epsilon_{1}&0&\cdots&0\\ 0&0&\cdots&0&0&-\epsilon_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0&0&0&\cdots&-\epsilon_{L}\end{pmatrix}$
+$\equiv\text{diag}(\epsilon_{\mu},-\epsilon_{\mu})=E_{\text{diag}}.$ (44)
+If we define the new Nambu fermion operators $\bm{\Phi}$ and $\bm{\Phi}^{\dagger}$ in such a way that
+$\bm{\Psi}=U\bm{\Phi},$ (45)
+we can write $\hat{H}$ in diagonal form
+$\hat{H}=\bm{\Psi}^{\dagger}H\bm{\Psi}=\bm{\Phi}^{\dagger}U^{\dagger}HU\bm{\Phi}=\bm{\Phi}^{\dagger}E_{\text{diag}}\bm{\Phi}.$ (46)
+Similarly to $\bm{\Psi}$, we can define new fermion operators $\hat{\gamma}$ such that
+$\bm{\Phi}=\begin{pmatrix}\hat{\gamma}\\ \hat{\gamma}^{\dagger}\end{pmatrix}=U^{\dagger}\bm{\Psi}=\begin{pmatrix}U^{\dagger}&V^{\dagger}\\ V^{T}&U^{T}\end{pmatrix}\begin{pmatrix}\hat{\mathbf{c}}\\ \hat{\mathbf{c}}^{\dagger}\end{pmatrix}$.
+(47)
+More explicitly, we can write:
+$\begin{cases}\hat{\gamma}_{\mu}=\sum_{j=1}^{L}\left(U^{*}_{j\mu}\hat{c}_{j}+V^{*}_{j\mu}\hat{c}^{\dagger}_{j}\right),\\ \hat{\gamma}^{\dagger}_{\mu}=\sum_{j=1}^{L}\left(V_{j\mu}\hat{c}_{j}+U_{j\mu}\hat{c}^{\dagger}_{j}\right),\end{cases}$ (48)
+which can be easily inverted, remembering that $\bm{\Psi}=U\bm{\Phi}$, to express the $\hat{c}_{j}$ operators in terms of the $\hat{\gamma}_{\mu}$:
+$\begin{cases}\hat{c}_{j}=\sum_{\mu}\left(U_{j\mu}\hat{\gamma}_{\mu}+V^{*}_{j\mu}\hat{\gamma}^{\dagger}_{\mu}\right),\\ \hat{c}^{\dagger}_{j}=\sum_{\mu}\left(V_{j\mu}\hat{\gamma}_{\mu}+U^{*}_{j\mu}\hat{\gamma}^{\dagger}_{\mu}\right).\end{cases}$ (49)
+Finally, $\hat{H}$ in terms of the $\hat{\gamma}$ operators reads, assuming we have taken $\epsilon_{\mu}>0$:
+$\hat{H}=\sum_{\mu=1}^{L}\left(\epsilon_{\mu}\hat{\gamma}^{\dagger}_{\mu}\hat{\gamma}_{\mu}-\epsilon_{\mu}\hat{\gamma}_{\mu}\hat{\gamma}^{\dagger}_{\mu}\right)=\sum_{\mu=1}^{L}2\epsilon_{\mu}\left(\hat{\gamma}^{\dagger}_{\mu}\hat{\gamma}_{\mu}-\frac{1}{2}\right),$ (50)
+Appendix C Adiabatic Transformation
 The time integral can be evaluated (assuming $t$ is large) as:
-[…]
+$\int_{0}^{t}e^{\frac{i}{\hbar}\Delta_{mn}t^{\prime}}\,dt^{\prime}\approx\frac{\hbar}{\Delta_{mn}}\left[e^{\frac{i}{\hbar}\Delta_{mn}t}-1\right]\approx\frac{2\hbar}{\Delta_{mn}}\sin\left(\frac{\Delta_{mn}t}{2\hbar}\right)$ (51)
 The transition amplitude becomes:
@@
 $P_{n\to m}(t)$ $=|a_{n\to m}(t)|^{2}\approx\left(\frac{2}{\Delta_{mn}}\right)^{2}\left|\langle m(\lambda)\middle|\frac{dH(\lambda)}{d\lambda}\middle|n(\lambda)\rangle\right|^{2}$
-×$\quad\times\left(\frac{d\lambda}{dt}\right)^{2}\sin^{2}\left(\frac{\Delta_{mn}t}{2\hbar}\right)$ (12)
+$\quad\times\left(\frac{d\lambda}{dt}\right)^{2}\sin^{2}\left(\frac{\Delta_{mn}t}{2\hbar}\right)$ (53)
 For small transitions, the oscillating term $\sin^{2}\left(\frac{\Delta_{mn}t}{2\hbar}\right)$, which oscillates between 0 and 1, can be ignored. Thus, the transition probability simplifies to:
@@
 This result forms the basis of the adiabatic condition, which states that for a process to be adiabatic, the rate of change $\frac{d\lambda}{dt}$ must be much smaller than the square of the energy difference $\Delta_{mn}^{2}$, thereby ensuring that the transition probability remains low.
-Appendix C Non perfect Thermalization
-In our paper, we worked on an ideal cycle with perfect ideal and adiabatic transformations, assuming the system is in contact with the two reservoirs for a time $t<a$, where $a$ is the thermalization time.
-According [18], the distribution becomes:
-$\langle b_{k}^{\dagger}b_{k}\rangle(\gamma)=f(\beta,\omega_{k})(1-e^{-J\gamma})+\langle b_{k}^{\dagger}b_{k}\rangle_{\rho_{0}}e^{-J\gamma},$ (15)
-We have two thermalization phases, from B to C and from D to A, each lasting $a/2$, and two adiabatic phases, from A to B and from C to D, each lasting $\gamma/2$.
-We assume $J=1$ here.
-Note that $J$ is a parameter that depends on the properties of the bath and the system-bath coupling, and obviously affects the thermalization time scale.
-The off-diagonal correlators and anomalous terms $\langle b_{k}b_{q}\rangle$ remain zero at any time, provided the initial condition for the Otto cycle is a thermal state (say, $\rho_{c}$).
-Let’s find an expression for the distribution after $n$ repetitions of the non-perfectly thermalized cycle. We introduce the diagonal $N\times N$ matrices :
-$\Theta_{c}(h)=\text{diag}\left(f\left[\beta_{c}(h),\omega_{k}\left(\lambda_{i}(f)\right)\right]\right)_{k=1,\ldots,N},$ (16)
-$\Gamma_{c}^{[n]}(h)=\text{diag}\left(\text{Tr}\left[b_{k}^{\dagger}b_{k}\rho_{c}^{[n]}(h)\right]\right)_{k=1,\ldots,N},$ (17)
-where $\rho_{c}^{[n]}(h)$ denotes the state at the end of the non-perfectly thermalized cycle with the bath at temperature $\beta_{c}^{-1}(h)$ after the $n$-th cycle repetition.
-At the end of the ideal adiabatic transformation, the system is put in contact with the reservoir at temperature $\beta^{-1}_{h}$ and reaches the state $\rho^{[1]}_{h}$, characterized by the distribution:
-$\Gamma^{[1]}_{h}=\Theta_{h}(1-e^{-\gamma})+\Gamma^{[0]}_{c}e^{-\gamma}.$ (18)
-After the ideal adiabatic transformation back, the system is put in contact with the reservoir at temperature $\beta^{-1}_{c}$ reaching the state $\rho^{[1]}_{c}$, characterized by the distribution:
-$\Gamma^{[1]}_{c}=\Theta_{c}(1-e^{-\gamma})+\Gamma^{[1]}_{h}e^{-\gamma}.$ (19)
-By imposing the stationary condition that the distributions $\Gamma$ are no longer dependent on the iteration, we obtain the following stationary solution:
-$\Gamma^{\infty}_{c}=h(\gamma)\left(\Theta_{c}+e^{-\gamma}\Theta_{h}\right),$ (20)
-$\Gamma^{\infty}_{h}=h(\gamma)\left(\Theta_{h}+e^{-\gamma}\Theta_{c}\right),$ (21)
-where $h(\gamma)=(1+e^{-\gamma})^{-1}$.
-The equations above suggest that, for the non-perfectly thermalized cycle ($n$-th), the heat absorbed and the work done per cycle are given by:
-$Q^{n\text{-th}}_{h}=g(\gamma)Q_{h},\quad W^{n\text{-th}}=g(\gamma)W,$ (22)
-where $g(\gamma)=(1-e^{-\gamma})h(\gamma)=\tanh(\gamma/2)$.
-The efficiency of the thermal engine in the case of the non-perfectly thermalized cycle remains the same as in the ideal case.
-For the case where we have a cycle with thermalization time and adiabatic time ($\tau=a/2$) and quench time ($T=b/2$):
-$P(a,b)=\frac{W^{a+b}}{a+b}=p(a,b)W,$ (23)
-where the function $p(a,b)$ is defined by:
-$p(a,b)\equiv\frac{\tanh(b/2)}{a+b}.$ (24)
-This relation shows that the power of the engine depends on the two parameters $a$ and $b$, which respectively characterize the quench time and the thermalization time.
-Appendix D Scaling Factor
-A high scaling factor ($\Pi$) indicates that, despite having an efficiency ($\eta$) lower than the Carnot efficiency, the engine is able to generate a significant amount of work.
-This suggests that the system is effectively optimized to extract as much work as possible, even under conditions far from the ideal efficiency.
-In quantum engines, a higher scaling factor reflects a more efficient conversion of thermal energy into work by leveraging quantum effects such as correlations and superposition.
-This implies that the system is able to utilize its quantum properties more effectively to enhance its thermodynamic performance.
-Furthermore, a high scaling factor can also indicate reduced energy losses, such as thermal dissipation or decoherence, allowing the engine to better harness the absorbed energy and maintain an efficiency closer to the Carnot limit.
-Thus, a higher scaling factor signifies a system that is not only more efficient but also more resilient to fluctuations and phase transitions, making better use of available energy resources to produce work, even if its absolute efficiency remains below that of Carnot.
-In summary, it reflects the system’s ability to maintain a favorable balance between the generated work and its deviation from the maximum theoretical efficiency.
-In conclusion, a high scaling factor is a global indicator of performance within a system, suggesting not only better energy conversion into work but also increased resilience to perturbations and more efficient use of available energy.
-This applies to both classical and quantum engines, where the effects of fluctuations and phase transitions become significant.
 ACKNOWLEDGMENTS
 A.H. acknowledges the financial support of the National Center for Scientific and Technical Research (CNRST) through the “PhD-Associate Scholarship-PASS” program.
+Declaration of Competing Interest:
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this manuscript.
+Data Availability:
+No data were used for the research described in this article.